### v3

# Identifiability of VAR(1) model in a stationary setting

Abstract
We consider a classical First-order Vector AutoRegressive (VAR(1)) model, where we interpret the autoregressive interaction matrix as influence relationships among the components of the VAR(1) process that can be encoded by a weighted directed graph. A majority of previous work studies the structural identifiability of the graph is based on time series observations and therefore relies on dynamical information. By constract, this work aims to incorporate the sampling insufficiency problem in ecological research, that is, due to limited resources, time-series observations are often unavailable and what we have in practice is independently, identically distributed data sampled from the stationary distribution, making classical time-series methods inapplicable. Within this framework, we assume that an equilibrium of VAR(1) exists, and study instead the identifiability of the graph from the stationary distribution, meaning that we seek a way to reconstruct the influence graph underlying the dynamic network using only static information. Importantly, the graph associated to the support of the autoregressive interaction matrix of the VAR(1) model does not coincide with the graphical model of the stationary distribution, making standard tools developped for Gaussian Graphical Models (GGMs) irrelevent. Therefore, this paper applies an approach from algebraic statistics that characterizes models using the Jacobian matroids associated with the parametrization of the models. We successfully derive easily applicable sufficient graphical conditions under which different graphs yield distinct steady-state distributions through the lens of maximal class, a concept that is initially introduced by this paper. Additionally, we illustrate how our results could be applied to characterize networks inspired by ecological research.
Keywords: model identifiability, vector autoregressive models (VAR), algebraic matroids, directed graph
MSC 2020: 62R01, 62H22, 62A09
1 Introduction
VAR(1) model.
This paper considers a classical First-order Vector AutoRegressive (VAR(1)) model:
[…]
where $\mathbf{x}_{t}$ and $\bm{\epsilon}_{t}$ are random vectors in $\mathbb{R}^{n}$ for some $n\in\mathbb{N}$. $\bm{\epsilon}_{t}$ is independent of $\mathbf{x}_{t-1}$, and $(\bm{\epsilon}_{t})_{t\in\mathbb{N}}$ are i.i.d.. $\Lambda=\left(\lambda_{ij}\right)\in M_{n}\left(\mathbb{R}\right)$, i.e. it is a deterministic $n\times n$ matrix with values in $\mathbb{R}$, which we call the interaction matrix. Each element $\lambda_{ij}$ represents the direct influence of the $i^{\mbox{th}}$ coordinate of $\mathbf{x}_{t-1}$ on the $j^{\mbox{th}}$ coordinate of $\mathbf{x}_{t}$. In this paper, the error term $\bm{\epsilon}_{t}$ is centered with covariance matrix $\omega I_{n}$, where $I_{n}\in M_{n}(\mathbb{R})$ is the identity matrix, and $\omega\in\mathbb{R}^{+}$ is a positive constant.
If the eigenvalues of $\Lambda$ are all smaller than $1$ in absolute value, then as $t$ goes to infinity, $x_{t}$ converges in distribution to a centered distribution with covariance matrix $\Sigma=\left(\sigma_{ij}\right)\in M_{n}(\mathbb{R})$ ([2]). $\Sigma$ satisfies ([21]):
$\Sigma=\Lambda^{T}\Sigma\Lambda+\omega I_{n}.$ (2)
Apply the vec operator ([15, Section 10.2]) on both sides, (2) is equivalent to:
$\mbox{vec}\left(\Sigma\right)=\left(I_{n}-\Lambda^{T}\otimes\Lambda^{T}\right)^{-1}\mbox{vec}\left(\omega I_{n}\right),$ (3)
where $\otimes$ is the matrix tensor product ([21]). (3) indicates that $\Sigma$ is unique given a unique set of parameters $\Lambda$ and $\omega$.
Problem statement.
This paper studies the identifiability of the VAR(1) model in a stationary setting, that is, whether different parameters ($\Lambda$ and $\omega$) produce the same $\Sigma$, using (2). We restrict the problem to the identifiability of the support of $\Lambda$, i.e. the location of its non-zero elements, which corresponds to the reconstruction of the underlying graph structure. Importantly, the properties of Gaussian Graphical Models (GGMs) (see [12] for an introduction) are inapplicable to our setting because the graphical model of the stationary distribution is the graph corresponding to the support of the inverse of the covariance matrix $\Sigma^{-1}$, and it is not the one induced by the support of $\Lambda$. For instance, Example 1 illustrates a case where $\Sigma^{-1}$, which is calculable from $\Sigma$, is complete (i.e. all elements are non-zero) while $\Lambda$ is not. Hence, it is necessary to seek for other methods to treat the problem.
To be more precise about the graphical model, we consider in this paper, the VAR(1) model (1) is associated with a directed graph $G=\left(V,\mathfrak{E}_{G}\right)$ with node set $V=\left\{1,2,\cdots,n\right\}$, each corresponding to one of the components of $\mathbf{x}_{t}$, and edge set $\mathfrak{E}_{G}$ consists of ordered pairs $\left(i,j\right)$ or $i\rightarrow j$, which represents the direct influence among components of $\mathbf{x}_{t}$. For any $i,j\in\left[n\right]$, an edge $\left(i,j\right)$ in $\mathfrak{E}_{G}$ indicates $\lambda_{ij}\neq 0$, meaning that there is a direct effect of the $i^{\mbox{th}}$ coordinate of $\mathbf{x}_{t-1}$ on the $j^{\mbox{th}}$ coordinate of $\mathbf{x}_{t}$. Figure 1 shows an example of a graphical VAR(1) model with $4$ nodes, and the corresponding graph $G$, where self-loops are omitted. Because $\Lambda$ is deterministic, i.e. it does not change over time, the dynamic model in $(a)$ is encoded by the graph $G$ in $(b)$ for simplicity. We assume throughout the paper that self-loops always exist in $G$, i.e. $(i,i)\in\mathfrak{E}_{G},\forall i\in\left[n\right]$, which is reasonable as it aligns with many real-life scenarios such as ecological networks where we expect the evolution of species to be dependent on itself.
(a) The graphical VAR(1) model
(b) The encoded graph $G$
Figure 1: Example of a graphical VAR(1) model and the encoded graph $G$
Motivation.
The motivation of this paper can be found in ecological research. Ecosystems are often modeled by dynamical systems. For example, in the model of Robert May ([14]), the linearization of the dynamics of abundancies or densities of $n$ interacting species around their equilibrium leads to the system:
$\frac{\mathrm{d}\mathbf{x}}{\mathrm{d}t}=\mathbf{Ax},$ (4)
where $\mathbf{A}$ is a deterministic $n\times n$ matrix that represents interactions among species through various mechanisms such as competition, predation or facilitation. The equation (1) can be seen as a stochastic and discrete in time version of (4).
VAR(1) models can also arise when we consider the (vectorial) fluctuation process $\mathbf{x}_{t}$ associated with individual-based models converging to Lotka-Volterra ordinary differential equations. The fluctuation process is then shown to satisfy a continuous time version of (1) with a noise having a diagonal covariance matrix (see [1] for more precisions). More importantly, it is reasonable to start from a basic and classical model also because, as pointed out later in this section, the study of dynamical system with only information from a steady state is a subject rarely explored.
The reason for which we are particularly considering a stationary setting lies in the sampling insufficiency problem of ecological research. It is a central topic for ecologists to reconstruct the dynamic interaction network, such as the support of $\mathbf{A}$ in (4) or the support of $\Lambda$ in (1), from the abundance data captured in the ecosystem, since studying the dynamic influence among species is crucial to understanding the evolution of the abundances of species and thus helps to prevent extinction. However, as pointed out by [13], fine-scale replication sampling such as time-series sampling is generally too costly and are often sacrificed in practice. Instead, the limited resources (manpower, time, money) are used to cover as much ground as possible during each sampling campaign. Consequently, a considerable amount of abundance data available covers a large number of locations but each of them is sampled only once, making classical time-series analysis inapplicable. For example, [4] inferred a global ocean cross-domain plankton co-occurrence network with one-time-samplings of $200$ stations. It is true that there are cases where each of the sampling sites is sampled multiple times, but the temporal gap between each sampling campaign is generally too large. For instance, [22] sampled for plant–aphid–ant interactions on $18$ forest islands once a year from 2020 to 2021. There are also researches that collect data regularly, e.g. monthly, for each sampling site, but the data points are not enough to conduct classical time-series analysis, such as [17], which collected plant and humming bird abundance over $19$ sites monthly but only over a little more than a year.
To resolve the sampling insufficiency problem and potentially make the planning of sampling campaigns easier for ecologists, we seek for a way to reconstruct the dynamic interaction network of an ecosystem with only i.i.d. samples of a final stationary distribution of the dynamical system, instead of time-series observations. As a first step of solving this problem, this paper assumes that we have the final state of the VAR(1) model, i.e. the matrix $\Sigma$ in (2), from which we try to recover the dynamic interaction network, i.e. the support of $\Lambda$. Although in practice, we will have i.i.d. observations from the stationary distribution and never the true $\Sigma$, it is worth verifying that the network is identifiable with the true $\Sigma$, because if not, it’s highly unlikely to be identifiable with real data.
Related work.
To the best of our knowledge, the identifiability of VAR(1) parameters in a stationary setting has rarely been explored in the existing literature. A majority of previous work that has dealt with the identifiability of the VAR(1) model, and achieved significant and systematic results, relied on observations that were sampled as time series (e.g. [7], [19]). This is reasonable because the parameters involve dynamic interactions among the components of the random process $\mathbf{x}_{t}$, making it natural to expect that it should only be inferred from dynamic data. Nevertheless, there is a literature that deals with the identifiability of the interaction matrix $\Lambda$ of VAR(1) in a stationary setting. [21] provides a necessary identifiability condition based on the number of parameters to estimate by counting the parameters. However, this condition only excludes a small subset of cases where the interaction matrix is not identifiable, while we are more interested in sufficient conditions for identifiability. Young’s condition also didn’t provide enough insight into the link between dynamical network and the stationary distribution it generates, resulting in limited interpretability.
Our contribution.
We address the problem using methods from algebraic statistics (see [18] for an introduction). These algebraic approaches have already been proven fruitful by [5] in the context of linear structural equation models (SEMs), for which sufficient graphical conditions for identifiability are derived. We apply the methods for testing the identifiability of [5] to our settings and set up a framework for identifiability of VAR(1) graphs in a stationary setting. We represent the set of stationary distributions corresponding to a graph $G$ by the set of covariance matrices $\Sigma$, denoted as $\mathbf{M}_{G}$, and study the naturally associated algebraic variety defined as the Zariski closure of $\mathbf{M}_{G}$. In algebraic statistics, a finite family of graphs are generically identifiable (formal definition in Section 2) if the intersection of any pair of models $\mathbf{M}_{G_{1}}$, $\mathbf{M}_{G_{2}}$ associated with graphs $G_{1}$, $G_{2}$ is a Lebesgue measure zero subset of any of the two models, which could be characterized through the Jacobian structures of the parametrization. By applying these results into our setting, we successfully derive sufficient graphical conditions for generic identifiability. Our results are based on a concept that we introduced in this paper: “maximal classes”, which can be interpreted as sets of nodes that receive information from a common ”source” in a directed graph (formal definition and relevant reconstruction algorithms are provided in later sections). We prove that graphs with different maximal classes generally yield different stationary distributions. Our graphical conditions for generic identifiability are not only easy to verify in practice, but also highly interpretable, as maximal classes serve as a bridge between dynamical network and the corresponding stationary distribution. We also provide several illustrations of how to apply our results to networks inspired by ecological research.
Outline.
The rest of this paper is organized as follows: Section 2 formally defines the statistical model and introduces the definitions and testing criteria for global and generic identifiability with an overview of the main results of this paper in the end. Section 3 presents results on the Jacobian structure of the parametrization of the model. Section 4 presents the concept of “maximal classes” for directed graphs and relevant reconstruction algorithms, which is the core of this study. Section 5 introduces two sufficient conditions of generic identifiability based on maximal classes, one of which is only valid for characterizing models with the same dimension, while the other one can be extended to models with any dimension. A result that allows us to calculate the dimension of the model for graphs without ”multi-edges” is also provided to make the two conditions applicable. Section 6 is an illustration of how our results could be applied in ecological research.
2 Identifiability of the VAR(1) model in a stationary setting
In this section, we start with introducing the parametrization and statistical model of our settings, then define global and generic identifiability, along with the identifiability criteria used throughout this paper. Finally, we provide an overview of our main results, i.e., sufficient conditions for generic identifiability.
The definition and criteria for identifiability are based on the concept of a generic point. We define a point in a set as generic if it does not lie within a proper (i.e., strictly included) algebraic subvariety of the set. In concrete terms, a generic point is any point in the set except for those that satisfy a specific non-trivial polynomial equation. Because such proper algebraic subvarieties have Lebesgue measure zero, another interpretation is that a parameter value sampled from a continuous probability distribution is generic with probability one.
2.1 Problem settings
Given a VAR(1) model (1) and the associated directed graph $G=\left(V,\mathfrak{E}_{G}\right)$, where $V=[n]=\{1,2,\cdots,n\}$, define the set of interaction matrices $M_{G}$ associated to a graph $G$ as:
$M_{G}:=\left\{\Lambda=\left(\lambda_{ij}\right)\in M_{n}\left(\mathbb{R}\right)\mid\rho(\Lambda)<1,\mbox{ and }\lambda_{ij}=0\mbox{ if }\left(i,j\right)\notin\mathfrak{E}_{G}\right\},$
where $\rho(\cdot)$ is the spectral radius, and the condition $\rho(\Lambda)<1$ ensures the existence of the equilibrium. The parameter space is then $M_{G}\times\mathbb{R}^{+}$, corresponding to elements of $\Lambda$ and $\omega$.
Next, we formally define the model.
Definition 1.
For the stationary VAR(1) model corresponding to a graph $G$, define $\mathbf{M}_{G}$ the following set of matrices:
$\mathbf{M}_{G}=\left\{\Sigma\mid\Sigma=\phi_{G}\left(\Lambda,\omega\right),\Lambda\in M_{G},\omega\in\mathbb{R}^{+}\right\},$
where
$\phi_{G}:M_{G}\times\mathbb{R}^{+}$ $\rightarrow M_{n}\left(\mathbb{R}\right)$
$\left(\Lambda,\omega\right)$ $\mapsto\Sigma,\mbox{ s.t. }\Sigma=\Lambda^{T}\Sigma\Lambda+\omega I_{n}$
is the parametrization map defined by (2). By abuse of notation, we call the set $\mathbf{M}_{G}$ the stationary VAR(1) model.
Remark 1 (Structural Properties).
We highlight two structural characteristics of the parameter space $M_{G}\times\mathbb{R}^{+}$ and the model $\mathbf{M}_{G}$:
1. Nested structure: The definition of $M_{G}$ allows for edge weights to vanish (i.e., $\lambda_{ij}=0$ for $(i,j)\in\mathfrak{E}_{G}$). Consequently, the parameter space of any subgraph $G^{\prime}\subset G$ is naturally embedded within $M_{G}$. This inclusion property extends to the models (i.e., $\mathbf{M}_{G^{\prime}}\subseteq\mathbf{M}_{G}$), ensuring a continuous transition between the model $\mathbf{M}_{G}$ and its sub-models.
2. Error structure: The covariance matrix of the error term in (1) is assumed to be diagonal with identical values $\omega$. While this assumption is central to our calculations, the underlying framework is generalizable to other diagonal matrices with distinct entries. However, the calculations do not extend to non-diagonal matrices. We adopted the identical diagonal assumption to minimize the number of parameters to estimate, allowing us to focus specifically on the reconstruction of $\Lambda$.
In this paper, we study the identifiability of a finite family of stationary VAR(1) models, i.e. assume that we are given a finite list of possible graphs, we try to see under what conditions they yield different stationary distributions, and in particular different stationary covariance relationships. This is particularly useful in ecological research where we could have prior information about the network, and were able to reduce the network to a finite list of candidates. Besides, as illustrated in Example 1 below, identifiability of the underlying graph from the covariance matrix $\Sigma$ in a conventional sense is not achieved because there exist two networks that yield the same stationary distribution.
For a finite family of stationary VAR(1) models, we characterize them by studying the dimension of the intersection of the images of the parametrization map, in the sense that two models are generically identifiable if they only intersect in a set with smaller dimension (detailed definitions of identifiability introduced in the next section). The dimension of the model $\mathbf{M}_{G}$ is defined as the rank of the (transpose of the) Jacobian matrix of the parametrization map evaluated at a generic point. This definition coincides with the standard dimension of $\mathbf{M}_{G}$ viewed as a manifold. Formally, this corresponds to the algebraic dimension with respect to the Zariski topology as an algebraic variety (see Appendix A for details).
The next section presents the definitions for global and generic identifiability.
2.2 Global and generic identifiability
Definition 2.
Given a family of stationary VAR(1) models $\left\{\mathbf{M}_{k}\right\}_{k=1}^{K}$ and associated graphs $\left\{G_{k}=\left(V,\mathfrak{E}_{G_{k}}\right)\right\}_{k=1}^{K}$, where $K\in\mathbb{N}$, then the model, i.e. the discrete parameter $k$ is globally identifiable if for any distinct pair $\left(k_{1},k_{2}\right)$ of values from $1$ to $K$, $\mathbf{M}_{k_{1}}\cap\mathbf{M}_{k_{2}}=\emptyset$.
In fact, global identifiability is a highly restrictive property of the models. There exist cases where models with different underlying graphs, i.e. different supports of $\Lambda$, yield the same $\Sigma$ (see Example 1).
Example 1.
Consider the case where $n=3$, and the following two interaction matrices, which represent different graphs:
$\Lambda_{1}=\left[\begin{array}[]{ccc}0.50&0.70&0.00\\ 0.00&0.90&0.00\\ 0.00&0.80&0.40\end{array}\right],\quad\Lambda_{2}=\left[\begin{array}[]{ccc}0.50&0.67&-0.01\\ 0.00&0.94&0.02\\ 0.00&0.00&0.38\end{array}\right]$
and if we let $\omega$ in both cases equals to $1$, then by (2), both models yield the following covariance matrix:
$\Sigma=\left[\begin{array}[]{ccc}1.33&0.85&0.00\\ 0.85&22.85&0.60\\ 0.00&0.60&1.19\end{array}\right].$
Therefore, we focus instead on a less restrictive definition of identifiability, called generic identifiability, which is widely used in the field of algebraic statistics. It allows the images of parametrization maps to intersect only in a set with Lebesgue measure zero.
Definition 3.
Let $\left\{\mathbf{M}_{k}\right\}_{k=1}^{K}$ be a family of stationary VAR(1) models as in Definition 2. Then the discrete parameter $k$ is generically identifiable if for any distinct pair $\left(k_{1},k_{2}\right)$ of values from $1$ to $K$,
${\operatorname{dim}}\left(\mathbf{M}_{k_{1}}\cap\mathbf{M}_{k_{2}}\right)<\min\left\{{\operatorname{dim}}\left(\mathbf{M}_{k_{1}}\right),{\operatorname{dim}}\left(\mathbf{M}_{k_{2}}\right)\right\}.$
The geometric interpretation of Definition 3 is that a family of models are generically identifiable if the intersection of any two models in the family is a Lebesgue measure zero subset of both models. But this definition also implies that in some circumstances, models with different dimensions are not generically identifiable, which is not desired. Instead, the following definition of generic identifiability is used.
Definition 4.
Let $\left\{\mathbf{M}_{k}\right\}_{k=1}^{K}$ be a family of stationary VAR(1) models as above. Then the discrete parameter $k$ is generically identifiable if for each pair $\left(k_{1},k_{2}\right)$ of values of $k$, where $k_{1}\neq k_{2}$,
${\operatorname{dim}}\left(\mathbf{M}_{k_{1}}\cap\mathbf{M}_{k_{2}}\right)<\max\left\{{\operatorname{dim}}\left(\mathbf{M}_{k_{1}}\right),{\operatorname{dim}}\left(\mathbf{M}_{k_{2}}\right)\right\}.$
Definition 4 states that the models in the family are generically identifiable from each other if the intersection of any two models in the family is a Lebesgue measure zero subset of the union of the models. This definition immediately implies that models with different dimensions are generically identifiable, as stated by the following proposition.
Proposition 1.
Let $\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ be two stationary VAR(1) models. If
${\operatorname{dim}}\left(\mathbf{M}_{1}\right)\neq{\operatorname{dim}}\left(\mathbf{M}_{2}\right),$
then these two models are generically identifiable.
Proposition 1 is one of the main tools that we used in this paper to test generic identifiability. Note this proposition does not imply that two models whose graphs are nested are generically identifiable, because as illustrated by the third row of Table 1 at the end of the paper, adding one edge to a graph might not change the dimension of the corresponding model. On the other hand, this proposition indicates that if we find a way to access the dimensions of the models, we can focus on the identifiability of models with the same dimension, in which case Definitions 3 and 4 are equivalent.
In fact, there exist many tools to test generic identifiability [18, Chapter 16]. This paper applies one of them that uses the Jacobian matroid derived from the parametrization of the model, developed by [5]. The following section provides an introduction to this method.
2.3 Identifiability criteria
Before introducing the tools to test generic identifiability, we present some necessary algebraic notions.
Definition 5.
Let $\phi$ be a $\mathcal{C}^{1}$ map from $\mathbb{R}^{d}$ to $\mathbb{R}^{p}$: $\phi\left(\theta_{1},\cdots,\theta_{d}\right)=\left(\phi_{1}\left(\bm{\theta}\right),\cdots,\phi_{p}\left(\bm{\theta}\right)\right)$. Then the Jacobian matrix of this map is:
$\mathbf{J}_{\phi}=\left(\frac{\partial\phi_{i}}{\partial\theta_{j}}\right),1\leq i\leq p,1\leq j\leq d.$
Next we define the Jacobian matrix for the stationary VAR(1) model, which is the transpose of the usual Jacobian matrix of the parametrization map.
Definition 6.
Let $\mathbf{M}_{G}$ be a stationary VAR(1) model, and $\phi_{G}$ in Definition 1 parametrizes $\mathbf{M}_{G}$. Then the Jacobian matrix of the model is
$\mathbf{J}_{G}=\left(\frac{\partial\phi_{j}}{\partial\theta_{i}}\right),1\leq i\leq E_{G}+1,1\leq j\leq\frac{n(1+n)}{2},$
where $E_{G}$ is the number of edges in $G$, $\phi_{j}$’s are the distinct entries of $\Sigma$, and $\theta_{i}$’s are the entries of $\Lambda$ and $\omega$.
Now we introduce the Jacobian matroid, the main object used to characterize models. To do so, we first introduce the linear matroid associated with a matrix. The general definition of a matroid is omitted here, as it is not necessary for the tools used in this paper. We refer the interested readers to [18] for further details.
Definition 7.
Let $A\in M_{m\times n}\left(\mathbb{R}\right)$ be a matrix, then the linear matroid of $A$ is the pair $\left\{E,\mathcal{I}\right\}$, where $E=\{1,\cdots,n\}$, representing the $n$ columns of $A$, and
$\mathcal{I}=\left\{S\subseteq E\mid A^{S}\mbox{ are linearly independent}\right\},$
where $A^{S}$ represents the set of columns of $A$ corresponding to the coordinates $S$.
Example 2.
Let
$A=\left[\begin{array}[]{cccc}2&0&2&4\\ 0&1&1&0\\ 0&0&1&0\\ 4&0&4&8\end{array}\right],$
then the linear matroid of $A$ is $\left\{E,\mathcal{I}\right\}$, where $E=\left\{1,2,3,4\right\}$, representing the $4$ columns of $A$ respectively, and
$\mathcal{I}=\left\{\emptyset,\left\{1\right\},\left\{2\right\},\left\{3\right\},\left\{4\right\},\left\{1,2\right\},\left\{1,3\right\},\left\{2,3\right\},\left\{2,4\right\},\left\{3,4\right\},\left\{1,2,3\right\},\left\{2,3,4\right\}\right\}.$
Note that $\left\{1,4\right\}$ is not included in the matroid because the first and the fourth column of $A$ are not independent, which also excludes $\left\{1,2,4\right\}$, $\left\{1,3,4\right\}$ and $\left\{1,2,3,4\right\}$.
Definition 8.
The Jacobian matroid of a stationary VAR(1) model $\mathbf{M}_{G}$, denoted as $\mathcal{J}\left(\mathbf{M}_{G}\right)$, is the linear matroid of the Jacobian matrix.
By abuse of notation, we identify $\mathcal{J}\left(\mathbf{M}_{G}\right)$ with $\mathcal{I}$ in Definition 7, representing the set of linearly independent columns sets of the Jacobian matrix.
Next, we introduce the criteria for generic identifiability based on Jacobian matroids. The criterion is extensively introduced in [18] and [10].
Proposition 2.
Let $\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ be two stationary VAR(1) models corresponding to graphs $G_{1}$ and $G_{2}$ with the same set of nodes. Without loss of generality, assume that ${\operatorname{dim}}\left({\mathbf{M}_{1}}\right)\geq{\operatorname{dim}}\left({\mathbf{M}_{2}}\right)$. If there exists a subset $S$ of the coordinates such that
$S\in\mathcal{J}\left(\mathbf{M}_{2}\right)\backslash\mathcal{J}\left(\mathbf{M}_{1}\right),$
then ${\operatorname{dim}}\left(\mathbf{M}_{1}\cap\mathbf{M}_{2}\right)<\min\left\{{\operatorname{dim}}\left(\mathbf{M}_{1}\right),{\operatorname{dim}}\left(\mathbf{M}_{2}\right)\right\}$.
Proof.
We refer the reader to [10, Proposition 10] for a proof of this result. ∎
In Section 5.2, we will prove that the dimension of a stationary VAR(1) model. i.e. the rank of the Jacobian matrix, is calculable for a subset of graphs. Therefore, we can focus specially on the identifiability of models with the same dimension, in which case the role of the two models are equivalent.
Proposition 3.
Let $\left\{\mathbf{M}_{k}\right\}_{k=1}^{K}$ be a finite set of stationary VAR(1) models corresponding to graphs $\left\{G_{k}\right\}_{i=1}^{K}$ with the same set of nodes and the same dimension. These models are generically identifiable, that is, the discrete parameter $i$ is generically identifiable, if for all $k_{1},k_{2}\in\left[K\right]$, where $k_{1}\neq k_{2}$, there exists a subset $S$ of the coordinates such that
[…]
Proof.
This is a direct result of Proposition 2. ∎
Propositions 2 and 3 will be the main tools used throughout this paper to test generic identifiability.
2.4 Main results
In this paper, we introduce a new concept for directed graphs named maximal classes (see Section 4). If we regard the directed graph as a network, a maximal class could be interpreted as a set of nodes that receive information from a common source. In a way, the set of maximal classes captures the dynamic information of the network and presents it in a static manner. We demonstrate that maximal classes are strongly linked to the Jacobian matroid of the model, based on which we conclude that for models with the same dimension, if they have different sets of maximal classes, they are generically identifiable (Theorem 3). Extending this result, for a general family of models (not necessarily with the same dimension), we then propose a more restrictive sufficient condition for generic identifiability based on maximal classes (Theorem 4). Finally, we propose a way to calculate the dimension of the model for a subset of graphs that do not contain multi-edges (Theorem 5), which enables us to characterize more models since different dimensions indicate generic identifiability by Definition 4. A summary and illustration of all the results in this paper are presented in Table 1 at the end of Section 5.
Together, these results provide a rigorous framework for addressing the issue of model identifiability with potential applicability to other parameterized models. Theorems 3 and 4 link model identifiability directly to the distinguishing of underlying graphs, offering a straightforward and practical identifiability criterion that is easy to apply. Moreover, the ability to calculate the rank of the Jacobian matrix shown in Theorem 5 provides valuable insight into the size of the space of possible probability distributions for a given set of parameters, which is itself already informative.
3 Jacobian Structure of VAR(1) in the Stationary Setting
In order to study the Jacobian matroid of the model, we need to calculate the Jacobian matrix. The goal of this section is to provide an explicit formula for the Jacobian matrix $\mathbf{J}_{G}$ for any directed graph $G$. Starting by deriving the Jacobian matrix for complete graphs, we then define a projection that maps it to that of any graphs, and in the end, we present a simplified formula for $\mathbf{J}_{G}$ based on the projection.
For any stationary VAR(1) model $\mathbf{M}_{G}$, the Jacobian matrix $\mathbf{J}_{G}$ is a matrix of size $\left(E_{G}+1\right)\times\left(n(n+1)\right)/2$, where $E_{G}=|\mathfrak{E}_{G}|$ is the number of edges in $G$. Each row of $\mathbf{J}_{G}$ corresponds to $\lambda_{ij}$ where $(i,j)\in\mathfrak{E}_{G}$ or $\omega$, and each column corresponds to $\sigma_{ab}$ where $a\leq b\in[n]$. Note that we only consider $a\leq b$ because $\sigma_{ab}=\sigma_{ba}$ for all $a,b\in[n]$.
Consider a stationary VAR(1) model $\mathbf{M}_{\overline{G}}$ with $\overline{G}=\left(V,\mathfrak{E}_{\overline{G}}\right)$ complete, i.e. $\mathfrak{E}_{\overline{G}}=\left\{(i,j)\mid i,j\in[n]\right\}$. Denote the Jacobian matrix of this model as $\mathbf{J}$, which is of size $\left(n^{2}+1\right)\times\left(n(n+1)\right)/2$. Define an ”extended” Jacobian matrix $\overline{\mathbf{J}}$, which is $\mathbf{J}$ with additional columns that correspond to $\sigma_{ab}$ where $a>b\in[n]$, meaning that $\overline{\mathbf{J}}$ is of size $\left(n^{2}+1\right)\times n^{2}$. The following lemma provides a formula for this $\overline{\mathbf{J}}$.
Lemma 1.
Let $\mathbf{M}_{G}$ be a stationary VAR(1) model with $G$ complete, and $P_{i,j}\in M_{n}\left(\mathbb{R}\right)$ for all $i,j\in[n]$ be the identity matrix switching the $i^{th}$ and $j^{th}$ columns. Then the extended Jacobian matrix $\overline{\mathbf{J}}$ satisfies:
$\overline{\mathbf{J}}=\overline{\mathbf{J}}\left(\Lambda\otimes\Lambda\right)+B,$
where
$B=\left[\begin{array}[]{c}\left(\Sigma\Lambda\otimes I_{n}\right)\left(I_{n^{2}}+\mathbf{P}\right)\\ \mbox{vec}\left(I_{n^{2}}\right)^{T}\end{array}\right],$
and
$\mathbf{P}=\prod_{i\geq 1}\prod_{j>i}P_{\left(i-1\right)n+j,\left(j-1\right)n+i}.$
Proof.
The proof is given in Appendix C. ∎
In Lemma 1, $P_{\left(i-1\right)n+j,\left(j-1\right)n+i}$ is the identify matrix exchanging the $i^{th}$ and $j^{th}$ columns of the $j^{th}$ and $i^{th}$ blocks, and hence $\mathbf{P}$ is just the indentity matrix exchanging columns n(n1)/2n(n-1)/2 times. The following example is a visualization of the function of $\mathbf{P}$.
Example 3.
When $n=3$, $\mathbf{P}$ does the following: for all $a,\cdots,i\in\mathbb{R}$,
$\left(\left[\begin{array}[]{ccc}a&b&c\\ d&e&f\\ g&h&i\end{array}\right]\otimes I_{3}\right)\mathbf{P}=\left[\begin{array}[]{ccccccccc}a&&&b&&&c&&\\ &a&&&b&&&c&\\ &&a&&&b&&&c\\ d&&&e&&&f&&\\ &d&&&e&&&f&\\ &&d&&&e&&&f\\ g&&&h&&&i&&\\ &g&&&h&&&i&\\ &&g&&&h&&&i\\ \end{array}\right]\mathbf{P}=\left[\begin{array}[]{ccccccccc}a&b&c&&&&&&\\ &&&a&b&c&&&\\ &&&&&&a&b&c\\ d&e&f&&&&&&\\ &&&d&e&f&&&\\ &&&&&&d&e&f\\ g&h&i&&&&&&\\ &&&g&h&i&&&\\ &&&&&&g&h&i\\ \end{array}\right]$
Let $\mathbf{M}_{G}$ be a stationary VAR(1) model with $G=\left(V,\mathfrak{E}_{G}\right)$, possibly not complete. Define again a corresponding ”extended” Jacobian matrix $\overline{\mathbf{J}}_{G}$, which is $\mathbf{J}_{G}$ with additional columns that correspond to $\sigma_{ab}$ where $a>b\in[n]$. Then, $\overline{\mathbf{J}}_{G}$ is of size $\left(E_{G}+1\right)\times n^{2}$. The following lemma proves that in fact, there is a simple relationship between $\overline{\mathbf{J}}$ and $\overline{\mathbf{J}}_{G}$, and we can define a projection from one to the other.
Lemma 2.
Let $\mathbf{M}_{G}$ be a stationary VAR(1) model with $G=\left(V,\mathfrak{E}_{G}\right)$ not necessarily complete, and $\overline{\mathbf{J}}$ be the extended Jacobian matrix for a complete graph. Define the projection
$\psi_{G}:M_{\left(n^{2}+1\right),n^{2}}\left(\mathbb{R}\right)$ $\rightarrow M_{\left(E_{G}+1\right),n^{2}}\left(\mathbb{R}\right)$
$\overline{\mathbf{J}}$ $\mapsto\psi_{G}\left(\overline{\mathbf{J}}\right),$
where $\psi_{G}$ removes the rows in $\overline{\mathbf{J}}$ that correspond to $\lambda_{ij}$, where $i,j\in\left[n\right]$ s.t. $\left(i,j\right)\notin\mathfrak{E}_{G}$, and sets such $\lambda_{ij}$ to zero in the remaining terms. Then
$\psi_{G}\left(\overline{\mathbf{J}}\right)=\overline{\mathbf{J}}_{G}.$
Sketch of proof.
The proof is based on the facts that $\overline{\mathbf{J}}_{G}$ has extra rows corresponding to $\lambda_{ij}$ where $(i,j)\notin\mathfrak{E}_{G}$, and that such $\lambda_{ij}$ equals zero in $\overline{\mathbf{J}}_{G}$. The complete proof is given in Appendix D. ∎
With Lemma 1 and Lemma 2, we now have a formula of $\overline{\mathbf{J}}_{G}$ for any directed graph $G$.
Lemma 3.
Let $\mathbf{M}_{G}$ be a stationary VAR(1) model. Then the extended Jacobian matrix satisfies:
$\overline{\mathbf{J}}_{G}=\overline{\mathbf{J}}_{G}\left(\Lambda\otimes\Lambda\right)+\psi_{G}(B),$
where $B$ is defined in Lemma 1.
Proof.
The proof is given in Appendix E. ∎
Here is an example of the extended Jacobian matrix.
Example 4.
Consider a stationary VAR(1) model $\mathbf{M}_{G}$ with a directed graph $G=\left(V,\mathfrak{E}_{G}\right)$, where $V=\left\{1,2,3\right\}$ and $\mathfrak{E}_{G}=\left\{\left(1,1\right),\left(2,2\right),\left(3,3\right),\left(1,3\right),\left(2,1\right)\right\}$. The graph and the corresponding interaction matrix are given below.
$\Lambda=\left[\begin{array}[]{ccc}\lambda_{11}&0&\lambda_{13}\\ \lambda_{21}&\lambda_{22}&0\\ 0&0&\lambda_{33}\end{array}\right].$
In this case, all computations can be explicit:
$\Sigma\Lambda=\left[\begin{array}[]{ccc}\sigma_{11}\lambda_{11}+\sigma_{12}\lambda_{21}&\sigma_{12}\lambda_{22}&\sigma_{11}\lambda_{13}+\sigma_{13}\lambda_{33}\\ \sigma_{21}\lambda_{11}+\sigma_{22}\lambda_{21}&\sigma_{22}\lambda_{22}&\sigma_{21}\lambda_{13}+\sigma_{23}\lambda_{33}\\ \sigma_{31}\lambda_{11}+\sigma_{32}\lambda_{21}&\sigma_{32}\lambda_{22}&\sigma_{31}\lambda_{13}+\sigma_{33}\lambda_{33}\end{array}\right].$
Then the extended Jacobian matrix $\overline{\mathbf{J}}_{G}$ satisfies:
$\overline{\mathbf{J}}_{G}=\overline{\mathbf{J}}_{G}A+\psi_{G}\left(B\right),$
where
$A$ $=\Lambda\otimes\Lambda$
$=\left[\begin{array}[]{ccc}\begin{array}[]{ccc}\lambda_{11}^{2}&0&\lambda_{11}\lambda_{13}\\ \lambda_{11}\lambda_{21}&\lambda_{11}\lambda_{22}&0\\ 0&0&\lambda_{11}\lambda_{33}\end{array}&\mathbf{0}&\begin{array}[]{ccc}\lambda_{11}\lambda_{13}&0&\lambda_{13}^{2}\\ \lambda_{13}\lambda_{21}&\lambda_{13}\lambda_{22}&0\\ 0&0&\lambda_{13}\lambda_{33}\end{array}\\ \begin{array}[]{ccc}\lambda_{21}\lambda_{11}&0&\lambda_{21}\lambda_{13}\\ \lambda_{21}^{2}&\lambda_{21}\lambda_{22}&0\\ 0&0&\lambda_{21}\lambda_{33}\end{array}&\begin{array}[]{ccc}\lambda_{22}\lambda_{11}&0&\lambda_{22}\lambda_{13}\\ \lambda_{22}\lambda_{21}&\lambda_{22}^{2}&0\\ 0&0&\lambda_{22}\lambda_{33}\end{array}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\begin{array}[]{ccc}\lambda_{33}\lambda_{11}&0&\lambda_{33}\lambda_{13}\\ \lambda_{33}\lambda_{21}&\lambda_{33}\lambda_{22}&0\\ 0&0&\lambda_{33}^{2}\end{array}\end{array}\right],$
and
$\psi_{G}\left(B\right)$ $=\left[\begin{array}[]{ccccc}2\left(\sigma_{11}\lambda_{11}+\sigma_{12}\lambda_{21}\right)&\sigma_{12}\lambda_{22}&\sigma_{11}\lambda_{13}+\sigma_{13}\lambda_{33}&\sigma_{12}\lambda_{22}&0\\ 0&0&\sigma_{11}\lambda_{11}+\sigma_{12}\lambda_{21}&0&0\\ 2\left(\sigma_{21}\lambda_{11}+\sigma_{22}\lambda_{21}\right)&\sigma_{22}\lambda_{22}&\sigma_{21}\lambda_{13}+\sigma_{23}\lambda_{33}&\sigma_{22}\lambda_{22}&0\\ 0&\sigma_{21}\lambda_{11}+\sigma_{22}\lambda_{21}&0&\sigma_{21}\lambda_{11}+\sigma_{22}\lambda_{21}&2\sigma_{22}\lambda_{22}\\ 0&0&\sigma_{31}\lambda_{11}+\sigma_{32}\lambda_{21}&0&0\\ 1&0&0&0&1\end{array}\right$.
$\left.\begin{array}[]{cccc}0&\sigma_{11}\lambda_{13}+\sigma_{13}\lambda_{33}&0&0\\ \sigma_{12}\lambda_{22}&\sigma_{11}\lambda_{11}+\sigma_{12}\lambda_{21}&\sigma_{12}\lambda_{22}&2\left(\sigma_{11}\lambda_{13}+\sigma_{13}\lambda_{33}\right)\\ 0&\sigma_{21}\lambda_{13}+\sigma_{23}\lambda_{33}&0&0\\ \sigma_{21}\lambda_{13}+\sigma_{23}\lambda_{33}&0&\sigma_{21}\lambda_{13}+\sigma_{23}\lambda_{33}&0\\ \sigma_{32}\lambda_{22}&\sigma_{31}\lambda_{11}+\sigma_{32}\lambda_{21}&\sigma_{32}\lambda_{22}&2\left(\sigma_{31}\lambda_{13}+\sigma_{33}\lambda_{33}\right)\\ 0&0&0&1\end{array}\right].$
In this example, if $\left(I_{n^{2}}-A\right)$ is invertible, then we will have an explicit formula for $\overline{\mathbf{J}}_{G}$.
From Lemma 3, it’s clear that if $\left(I_{n^{2}}-\Lambda\otimes\Lambda\right)$ is invertible, we would be able to define $\overline{\mathbf{J}}_{G}$ as a function of $\Lambda$ and $\omega$. The following proposition states that the set of the parameters such that the matrix $\left(I_{n^{2}}-\Lambda\otimes\Lambda\right)$ is not invertible has Lebesgue measure zero. Because of the generic settings, we never work with any set of parameters that has Lebesgue measure zero. Hence, we can assume that $\left(I_{n^{2}}-\Lambda\otimes\Lambda\right)$ is invertible.
Proposition 4.
Let $\mathbf{M}_{G}$ be a stationary VAR(1) model, $\mu_{G}$ be a measure defined on $M_{G}$, which is the Lebesgue measure on $\mathbb{R}^{E_{G}}$, and $M_{G}^{0}$ be the subset of $M_{G}$ defined as
$M_{G}^{0}:=\left\{\Lambda\in M_{G}\mid\left(I_{n^{2}}-\Lambda\otimes\Lambda\right)\mbox{ is not invertible}\right\},$
then
$\mu_{G}\left(M_{G}^{0}\right)=0,$
Proof.
It’s clear that $\left(I_{n^{2}}-\Lambda\otimes\Lambda\right)$ is not invertible if and only if $\det\left(I_{n^{2}}-\Lambda\otimes\Lambda\right)=0$. Since the determinant is a result of summations and multiplications of the elements of the matrix, it is a polynomial with respect to the entries of $\Lambda$, i.e. $\det\left(I_{n^{2}}-\Lambda\otimes\Lambda\right)\in\mathbb{R}\left[\lambda_{ij}\mid\left(i,j\right)\in\mathfrak{E}_{G}\right]$.
Let $\Lambda_{0}=2I_{n}$, then $\Lambda_{0}\in M_{G}$ for any graph $G$, and
$\det\left(I_{n^{2}}-\Lambda_{0}\otimes\Lambda_{0}\right)=\det\left(-3I_{n^{2}}\right)=\left(-3\right)^{n^{2}}\neq 0.$
Since $\det\left(I_{n^{2}}-\Lambda\otimes\Lambda\right)$ is a polynomial, by continuity, it does not identically equal to 0. Therefore, the set of $\Lambda$ such that the determinant equals zero must have zero measure, i.e. $\mu_{G}\left(M_{G}^{0}\right)=0$. ∎
The idea of the proof of Proposition 4 is original and will be applied several times in the following sections. Note that in Definition 3, we also used the Lebesgue measure, but defined on $\mathbb{R}^{\frac{n(1+n)}{2}}$. Here, we only have $E_{G}$ values of $\Lambda$ that vary, so $\mu_{G}$ is the Lebesgue measure defined on $\mathbb{R}^{E_{G}}$.
With Proposition 4, we now have a formula for the extended Jacobian matrix.
Theorem 1.
Let $\mathbf{M}_{G}$ be a stationary VAR(1) model. Then the extended Jacobian matrix is generically:
$\overline{\mathbf{J}}_{G}=\psi_{G}(B)\left(I_{n^{2}}-\Lambda\otimes\Lambda\right)^{-1},$
where $B$ is defined in Lemma 1.
Proof.
This is a direct result of Lemma 3 and Proposition 4. ∎
Theorem 1 enables us to calculate the extended Jacobian matrix $\overline{\mathbf{J}}_{G}$ for any directed graph $G$, meaning that we can also calculate the Jacobian matrix $\mathbf{J}_{G}$, since $\overline{\mathbf{J}}_{G}$ is just $\mathbf{J}_{G}$ with additional columns that are identical to a subset of columns of $\mathbf{J}_{G}$. Recall that the Jacobian matroid is the set of coordinates of columns that are linearly independent. The next section presents a new concept called ”maximal classes”, for any directed graph, which is highly related to the linear independence of the Jacobian matrix, and is at the core of this study.
4 Maximal classes
In this section, we introduce the definition and properties of maximal classes and highlight their relationship with some parameters of our model. In the following sections, we will see that in general, different maximal classes could imply generic identifiability. Section 4.1 formally defines maximal classes for directed graphs, and proves the uniqueness of the set of maximal classes for a given graph. Additionally, we propose Algorithm 1 to identify the set of maximal classes from any graph, and Algorithm 2 to reconstruct the list of graphs associated with a given set of maximal classes. We also compare the concept of maximal classes with related notions in graph theory. In Section 4.2, we highlight the strong relationship between maximal classes and the covariance matrix of the stationary distribution $\Sigma$, based on which another algorithm (Algorithm 3) is designed to reconstruct the set of maximal classes from the support of $\Sigma$. With Algorithm 2, the list of graphs associated with $\Sigma$ can also be reconstructed. Algorithms 3 and 2 serve as a first step to the reconstruction of the graph corresponding to a stationary VAR(1) model, which even though is not the main subject of this paper, still could inspire future works.
4.1 Definition and properties
Recall that a Strongly Connected Component (SCC) of a directed graph is a subgraph where each node can be reached by a directed path from any other node in the subgraph. As this defines an equivalence relation, a directed graph can be uniquely decomposed into its strongly connected components. By abuse of notation, we name an SCC a set of nodes that belong to the corresponding subgraph.
We know that the in-degree of a node $v$ in a directed graph is the number of edges directed to $v$. Similarly, we define the in-degree for a SCC, which is the number of edges coming from a node outside of the SCC and pointing to one of the nodes in the SCC. In other words, let $G=(V,\mathfrak{E}_{G})$ be a directed graph containing a SCC: $\left\{v_{1},\cdots,v_{p}\right\}$ for some $p\in\mathbb{N}$, the in-degree of the SCC is:
[…]
From now on, the in-degree of a node or a SCC in a directed graph $G$ means the in-degree excluding self-loops.
A node or a SCC in a directed graph with in-degree zero is special because when the graph is viewed as a network, it can be thought of as ”feeding” the network. Such node or SCC is defined as a source of the graph as edges are only directed away from it. Moreover, we define any node that is itself a source or belongs to a SCC that is a source as a source node of the graph. A maximal class is the set of all nodes that receive information from a given source node, which is formally defined below:
Definition 9.
Let $i$ be a source node of a directed graph. The maximal class $\mathcal{MC}_{[i]}$ is a set of nodes such that there exists a directed path from a source node $i$ to any node in the set and that is maximal with respect to inclusion, i.e. the set of all reachable nodes from a source node $i$. $i$ is defined as the source node of the maximal class. The source node $i$ or the SCC containing $i$ that is a source is defined as the source of the maximal class.
By definition, a maximal class might have multiple source nodes, but it can have only one SCC source.
The following lemma provides a necessary and sufficient condition for any two nodes to belong to the same maximal class. It can also serve as an equivalent definition of maximal classes, which will be useful in the following sections.
Lemma 4.
Let $G=\left(V,\mathfrak{E}_{G}\right)$ be a directed graph. Then two nodes $i,j\in V$ belong to the same maximal class of $G$ if and only if one of the following conditions is satisfied:
1. There exists a directed path in $G$ between $i$ and $j$.
2. There exists $l\neq i,j\in V$ such that there are directed paths from $l$ to $i$ and $j$ respectively.
Proof.
The proof is given in F. ∎
The following proposition proves the uniqueness of the set of maximal classes for any graph, which is a fundamental property of maximal classes.
Proposition 5.
Let $G$ be any directed graph, then the set of maximal classes associated with $G$ is unique.
Proof.
First, two maximal classes of a graph cannot have the same source. This is because by definition 9, a maximal class is the set of all nodes reachable by the source, and if two maximal classes have the same source, then they must be the same maximal class. Moreover, for any directed graph $G$, the set of sources is unique since the in-degree of any node or SCC is uniquely defined. Therefore, each source generates a unique maximal class. Hence, the set of maximal classes is unique. ∎
Example 5.
Figure 2: Example of a directed graph
The SCCs of the graph in Figure 2 are:
$\left\{1,2,3\right\},\{4\},\{5\},\{6\},$
It’s clear that $\left\{1,2,3\right\}$ and $\{5\}$ have in-degree zero, and hence are the sources. Therefore, the set of maximal classes is:
$\left\{\left\{\mathbf{1},\mathbf{2},\mathbf{3},4,6\right\},\left\{4,\mathbf{5},6\right\}\right\},$
where the nodes in bold are source nodes of the respective maximal classes. In other words, there are two maximal classes in this graph:
$\mathcal{MC}_{[1]}=\mathcal{MC}_{[2]}=\mathcal{MC}_{[3]}=\{1,2,3,4,6\}\mbox{, and }\mathcal{MC}_{[5]}=\{4,5,6\}.$
Because the set of maximal classes is unique (Proposition 5), we can design Algorithm 1 to derive the set of maximal classes for any directed graph. This Algorithm first applies Kosaraju’s algorithm ([16]) to find all SCCs of the graph, and then construct the condensed graph, which is a directed acyclic graph where all SCCs are represented as a single node. Finally, we perform a Depth First Search (DFS) algorithm ([20]) on each of the nodes with in-degree zero in the condensed graph.
The computational complexity of Algorithm 1 is dictated by the Kosaraju’s algorithm, the construction of the condensed graph, and the subsequent DFS searches. Since these components are all linear operations relative to the size of the graph, the complexity of the algorithm is upper bounded by $O(k(n+E))$, where $n$ is the number of nodes, $E$ is the number of edges in the original graph, and $k$ is the number of SCCs in the condensed graph. Consequently, the best-case scenario achieves a complexity of $O(n+E)$, while the worst-case upper bound is $O(n(n+E))$.
Algorithm 1 Algorithm to identify maximal classes of a directed graph
Input: a directed graph $G=\left(V,\mathfrak{E}_{G}\right)$
   Output: the number and the list of maximal classes of $G$
1:procedure DFS($G,i,j$) $\triangleright$ Find the set of all reachable nodes from $i$, and store it into the $j^{th}$ list of the list maxc
2:  add $i$ to maxc[$j$] $\triangleright$ Add $i$ itself to the list
3:  for all nodes $w\in V$ s.t. $\left(i,w\right)\in\mathfrak{E}_{G}$ do $\triangleright$ Do the same thing to all children of $i$
4:    DFS(G,w,j)(G,w,j)
5:procedure MAXC(G) $\triangleright$ Search for the maximal classes of graph $G$
6:  Perform Kosaraju’s algorithm, return $L$, a list of SCCs
7:  Construct the condensed graph $G^{\prime}=\left(L,\mathfrak{E}_{G^{\prime}}\right)$
8:  $j\leftarrow 1$
[…]
10:  for all $i\in|L|$ s.t. ${\operatorname{deg}}_{+}\left(L[i]\right)=0$ in $G^{\prime}$ do
[…]
12:    $j\leftarrow j+1$   return j1j-1 (number of maximal classes) and maxc (list of maximal classes)
To reconstruct the list of graphs associated with a given set of maximal classes, that is, the inverse problem of the precedent algorithm, we propose Algorithm 2. This algorithm starts with identifying candidate sources for each maximal class. It then iterates through the complete set of possible edges $\left(u,v\right)\in V\times V$ to isolate ’forbidden’ edges — those whose inclusion would violate the definitions of the maximal classes or their sources. Finally, the algorithm evaluates subsets of the remaining ’allowable’ edges, applying Algorithm 1 to verify which resulting graphs yield the target set of maximal classes.
The complexity of Algorithm 2 is significantly higher due to its combinatorial nature. While the initial source allocation and edge pruning is efficient at $O(m\cdot n^{2})$ with $m$ being the number of maximal classes, the subsequent exploration of edge subsets results in a worst-case complexity of $O(2^{p}\cdot n(n+E))$, where $p$ is the number of ’allowable’ edges. The resulting complexity is upper bounded by $O(m\cdot n^{2}+2^{p}\cdot n(n+E))$.
An R implementation of both the algorithms is available at https://github.com/Bi-xuan/maximal_class. While the procedures presented in this section provide a systematic way to reconstruct maximal classes from their corresponding graph structures as well as the inverse, it is important to note that these implementations are not computationally optimal and leave significant room for algorithmic refinement. However, such optimizations and a systematic empirical validation across a broad selection of synthetic data and large-scale simulated ecological networks remain outside the primary theoretical focus of this work. These enhancements and performance testing are dedicated to future research.
Algorithm 2 Algorithm for the reconstruction of graphs via maximal classes
Input: a list of maximal classes $\mathfrak{MC}=\left\{\mathcal{MC}_{k},k=1,\cdots,m\right\}$
   Output: the list of graphs associated with $\mathfrak{MC}$
1:procedure RECONSTRUCT($\mathfrak{MC}$)
2:  $V\leftarrow\bigcup_{k=1}^{m}\mathcal{MC}_{k}$ $\triangleright$ Define the set of nodes
3:  SourceCandidates $\leftarrow$ an empty list of lists
4:  for all $\mathcal{MC}_{k}\in\mathfrak{MC}$ do
5:    $S_{k}\leftarrow\left\{v\in\mathcal{MC}_{k}\mid v\notin\mathcal{MC}_{j},\forall j\neq k\right\}$ $\triangleright$ Identify nodes that belong to only one maximal class
6:    Add $S_{k}$ to SourceCandidates
7:  ForbiddenEdges $\leftarrow\emptyset$
8:  for all $(u,v)\in V\times V$ do
9:    for all $\mathcal{MC}_{k}\in\mathfrak{MC}$ do
10:     if $u\in\mathcal{MC}_{k}$ and $v\notin\mathcal{MC}_{k}$ then
11:      Add $(u,v)$ to ForbiddenEdges $\triangleright$ By definition of maximal classes
12:     if $v\in S_{k}$ and $u\notin S_{k}$ then
13:      Add $(u,v)$ to ForbiddenEdges $\triangleright$ Source nodes must have in-degree zero from outside of their SCC
14:  PossibleEdges $\leftarrow V\times V\backslashForbiddenEdges$
15:  PossibleGraphs $\leftarrow\emptyset$
16:  for all selection of edges $E\subseteq$ PossibleEdges do $\triangleright$ Permute over subsets of possible edges
17:    $G\leftarrow(V,E)$
18:    if MAXC($G$) $\mathfrak{MC}$ then $\triangleright$ Verify using Algorithm 1
19:     Add $G$ to PossibleGraphs       return PossibleGraphs
In fact, maximal classes are very similar to some common terminologies in graph theory. Recall that given a subset of nodes, an induced graph of the set of nodes is the subgraph that contains all edges in the original graph among the nodes within the set, a directed tree is a directed graph whose underlying undirected graph is a tree, a rooted tree is a directed tree with a designated node called the root and each edge is considered to be directed away from the root, a spanning tree of a graph is a spanning subgraph that is a tree, and a rooted spanning tree of a graph is a spanning tree that is rooted. This similarity is highlighted by the following proposition.
Proposition 6.
The induced subgraph of a maximal class of a directed graph $G$ has a rooted spanning tree, where the root is (one of) the source node(s) of the maximal class.
Proof.
For an induced subgraph of any maximal class of a directed graph $G$, keeping only the edges that direct away from (one of) the source node(s) of the maximal class would result in a rooted spanning tree. ∎
Note that the induced subgraphs of maximal classes do not construct a spanning forest because there might exist overlaps between different trees (Example 5).
4.2 Derivation of maximal classes from $\Sigma$
In this section, we describe the relationship between maximal classes and the covariance matrix $\Sigma$ and show that it is possible to reconstruct the set of maximal classes from the support of $\Sigma$.
We start with a Lemma on the values of the parameter $\Lambda$, which will be helpful in the subsequent calculations.
Lemma 5.
Let $\mathbf{M}_{G}$ be a stationary VAR(1) model, and $\lambda_{ij}^{[k]}=\left(\Lambda^{k}\right)_{ij}$, for all $k\in\mathbb{N}$, then generically
$\Lambda\in M^{P}\cap M^{S},$
where
[…]
$M^{S}:=$ $\left\{\Lambda\in M_{n}\left(\mathbb{R}\right)\mid\forall i,j\in V,\sum_{k=0}^{+\infty}\sum_{a=1}^{n}\lambda_{ai}^{\left[k\right]}\lambda_{aj}^{\left[k\right]}=0\Leftrightarrow\forall k\in\mathbb{N},a\in\left[n\right],\lambda_{ai}^{\left[k\right]}\lambda_{aj}^{\left[k\right]}=0,\right$.
$\left.\mbox{and }\sum_{s=1}^{n}\sum_{k=0}^{+\infty}\sum_{l=1}^{n}\lambda_{li}^{\left[k\right]}\lambda_{ls}^{\left[k\right]}\lambda_{sj}=0\Leftrightarrow\forall k\in\mathbb{N},s,l\in\left[n\right],\lambda_{li}^{\left[k\right]}\lambda_{ls}^{\left[k\right]}\lambda_{sj}=0\right\}.$
Proof.
The proof is given in Appendix G. ∎
In fact, $M^{P}$ is a set of matrices $\Lambda$ such that for all $i,j\in[n]$, if there exists a directed path from $i$ to $j$, then after a certain power, the corresponding coefficient $\lambda_{ij}$ stays non-zero. $M^{S}$ is a set of matrices $\Lambda$ such that the infinite sum of certain elements equals zero is equivalent to requiring that each part of the sum equals zero individually. In other words, we avoid cancellations among elements of different powers of $\Lambda$ in this set.
In this paper, because we consider generic identifiability, and hence generic properties of the parameters, we assume that for all stationary VAR(1) models, the associated interaction matrices satisfy: $\Lambda\in M^{S}\cap M^{P}$. It is reasonable because Lemma 5 implies that $M_{G}\backslash\left(M^{S}\cap M^{P}\right)$ has Lebesgue measure zero. Note that this type of argument is already mentioned in Proposition 4. This assumption is necessary for proving the relationships between maximal classes and the parameters, which are highlighted by Proposition 7 and Lemma 6.
Proposition 7.
Let $\mathbf{M}_{G}$ be a stationary VAR(1) model with corresponding directed graph $G=\left(V,\mathfrak{E}_{G}\right)$, then for any two nodes $i,j\in[n]$, $\sigma_{ij}=\sigma_{ji}=0$ if and only if $i$ and $j$ do not belong to the same maximal class of $G$.
Proof.
For all $s,t\in\left\{1,\cdots,n\right\}$, let $\lambda_{st}^{\left[k\right]}=\left(\Lambda^{k}\right)_{st}$. Assume $i$ and $j$ do not belong to the same maximal class, i.e. by Lemma 4, there is no directed path between $i$ and $j$, and additionally, they don’t have the same ”ancestor”. Therefore, by the properties of matrix multiplication, for all $k\in\mathbb{N}$,
$\lambda_{ij}^{\left[k\right]}=\lambda_{ji}^{\left[k\right]}=0,$
and for all $l\neq i,j\in\left[n\right]$,
$\lambda_{li}^{\left[k\right]}=0\mbox{ or }\lambda_{lj}^{\left[k\right]}=0.$
Proposition 4 states that $\left(I_{n^{2}}-\Lambda\otimes\Lambda\right)$ is generically invertible, which implies that $\left(I_{n^{2}}-\Lambda^{T}\otimes\Lambda^{T}\right)$ is generically invertible since $\left(I_{n^{2}}-\Lambda^{T}\otimes\Lambda^{T}\right)$ is just the transpose of $\left(I_{n^{2}}-\Lambda\otimes\Lambda\right)$. Hence, from (3),
$\mbox{vec}\left(\Sigma\right)$ $=\left(I_{n^{2}}-\Lambda^{T}\otimes\Lambda^{T}\right)^{-1}\mbox{vec}\left(\omega I_{n}\right)=\omega\left(\sum_{k=0}^{+\infty}\left(\Lambda^{T}\otimes\Lambda^{T}\right)^{k}\right)\mbox{vec}\left(I_{n}\right)$
$=\omega\left(\sum_{k=0}^{+\infty}\left(\Lambda^{T}\right)^{k}\otimes\left(\Lambda^{T}\right)^{k}\right)\mbox{vec}\left(I_{n}\right)=\omega\left(\sum_{k=0}^{+\infty}\left(\Lambda^{k}\right)^{T}\otimes\left(\Lambda^{k}\right)^{T}\right)\mbox{vec}\left(I_{n}\right)$
$=\omega\sum_{k=0}^{+\infty}\left[\begin{array}[]{ccccccc}\lambda_{11}^{\left[k\right]}\left(\Lambda^{k}\right)^{T}&\cdots&\lambda_{i1}^{\left[k\right]}\left(\Lambda^{k}\right)^{T}&\cdots&\lambda_{j1}^{\left[k\right]}\left(\Lambda^{k}\right)^{T}&\cdots&\lambda_{n1}^{\left[k\right]}\left(\Lambda^{k}\right)^{T}\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots\\ \lambda_{1i}^{\left[k\right]}\left(\Lambda^{k}\right)^{T}&\cdots&\lambda_{ii}^{\left[k\right]}\left(\Lambda^{k}\right)^{T}&\cdots&0&\cdots&\lambda_{ni}^{\left[k\right]}\left(\Lambda^{k}\right)^{T}\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots\\ \lambda_{1j}^{\left[k\right]}\left(\Lambda^{k}\right)^{T}&\cdots&0&\cdots&\lambda_{jj}^{\left[k\right]}\left(\Lambda^{k}\right)^{T}&\cdots&\lambda_{nj}^{\left[k\right]}\left(\Lambda^{k}\right)^{T}\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots\\ \lambda_{1n}^{\left[k\right]}\left(\Lambda^{k}\right)^{T}&\cdots&\lambda_{in}^{\left[k\right]}\left(\Lambda^{k}\right)^{T}&\cdots&\lambda_{jn}^{\left[k\right]}\left(\Lambda^{k}\right)^{T}&\cdots&\lambda_{nn}^{\left[k\right]}\left(\Lambda^{k}\right)^{T}\\ \end{array}\right]\left[\begin{array}[]{c}1\\ 0\\ \vdots\\ 0\\ 1\\ \vdots\\ 0\\ \vdots\\ 1\end{array}\right],$
where the second equality is true because the $\rho(\Lambda)<1$ implies $\rho(\Lambda^{T}\otimes\Lambda^{T})<1$. Therefore,
$\sigma_{ij}=\omega\sum_{k=0}^{+\infty}\sum_{l=1}^{n}\lambda_{li}^{\left[k\right]}\lambda_{lj}^{\left[k\right]}=0=\sigma_{ji},$
On the other hand, by Lemma 5, if for $i,j\in V$, $\sigma_{ij}=\sigma_{ji}=0$, then generically $\lambda_{li}^{\left[k\right]}\lambda_{lj}^{\left[k\right]}=0$ for all $k\in\mathbb{N}$ and $l\in V$. Therefore, there is no directed path between $i$ and $j$, and they do not have the same ancestor. Thus, they do not belong to the same maximal class. ∎
From Proposition 7, we know that there is a direct link between the support of $\Sigma$ and the set of maximal classes. Using this result, we are able to derive the set of maximal classes of the underlying graph from $\Sigma$.
For all $i\in[n]$, define the set of nodes that belong to the same maximal class as $i$:
$\mathcal{C}_{i}:=\{j\mid j\in[n],i,j\mbox{ belong to the same maximal class}\}.$
And from Proposition 7,
$\mathcal{C}_{i}=\{j\mid j\in[n],\sigma_{ij}\neq 0\}.$
Note that for all $i\in[n]$, $\mathcal{C}_{i}$ exists, but may not be a maximal class. For example, in Figure 2, $\mathcal{C}_{4}=\{1,2,3,4,5,6\}$ is not a maximal class, while $\mathcal{C}_{5}=\{4,5,6\}$ is a maximal class.
The following theorem proves that the set of maximal classes is the set of $\mathcal{C}_{i}$ excluding the ones containing two nodes that do not belong to the same maximal class.
Theorem 2.
Let $\mathbf{M}_{G}$ be a stationary VAR(1) model. Then the set of maximal classes $\mathfrak{MC}$ satisfies
$\mathfrak{MC}=\mathfrak{C}\backslash\mathfrak{C}^{\prime}$,
where
$\mathfrak{C}$ $=\left\{\mathcal{C}_{i}\mid i\in[n]\right\},$
$\mathfrak{C}^{\prime}$ $=\left\{\mathcal{C}_{i}\in\mathfrak{C}\mid i\in[n],\mbox{ and }\exists k,l\in\mathcal{C}_{i}\ s.t.\ \sigma_{kl}=0\right\}.$
Proof.
First, we prove that $\mathfrak{MC}\subseteq\mathfrak{C}$. Consider $\mathcal{MC}\in\mathfrak{MC}$, there exists a source node $i\in[n]$, s.t. $\mathcal{MC}=\mathcal{MC}_{[i]}$. By Proposition 7, for all $j\in\mathcal{MC}_{[i]},\sigma_{ij}\neq 0$, hence $j\in\mathcal{C}_{i}$. Therefore,
$\mathcal{MC}_{[i]}\subseteq\mathcal{C}_{i}.$
Assume $\exists j\in\mathcal{C}_{i}\backslash\mathcal{MC}_{[i]}$, then one of the following circumstances is true:
1. There exists a directed path from $i$ to $j$, then $j\in\mathcal{MC}_{[i]}$, which is a contradiction.
2. There exists a directed path from $j$ to $i$, then ${\operatorname{deg}}_{+}(i)\neq 0$, which means that the source of $\mathcal{MC}_{[i]}$ is a SCC containing $i$. Hence either $j$ belongs to the SCC, then $j\in\mathcal{MC}_{[i]}$, which is a contradiction, or $j$ does not belong to the SCC, then the in-degree of the SCC is non-zero, i.e. the SCC containing $i$ is not the source of $\mathcal{MC}_{[i]}$, which is also a contradiction.
3. There exists another node $c\in[n]$ s.t. there are paths from $c$ to $i$, and from $c$ to $j$ respectively, then ${\operatorname{deg}}_{+}(i)\neq 0$, which means that the source of $\mathcal{MC}_{[i]}$ is a SCC containing $i$. Hence either $c$ belongs to the SCC, then $j\in\mathcal{MC}_{[i]}$, which is a contradiction, or $c$ does not belong to the SCC, then the in-degree of the SCC is non-zero, i.e. the SCC containing $i$ is not the source of $\mathcal{MC}_{[i]}$, which is also a contradiction.
Therefore
$\mathcal{MC}_{[i]}=\mathcal{C}_{i}.$
This proves that $\mathfrak{MC}\subseteq\mathfrak{C}$.
On the other hand, consider $\mathcal{C}_{i}\in\mathfrak{C}^{\prime}$ such that $\mathcal{C}_{i}\notin\mathfrak{MC}$. Then $\mathcal{C}_{i}$ is not a maximal class, because it contains two nodes that do not belong to the same maximal class by definition. Therefore
$\mathfrak{MC}\subseteq\mathfrak{C}\backslash\mathfrak{C}^{\prime}$.
Let $\mathcal{C}_{i}\in\mathfrak{C}\backslash\mathfrak{C}^{\prime}$ for certain $i\in[n]$ and assume that $\mathcal{C}_{i}\notin\mathfrak{MC}$ (proof in Appendix F). We know that $i$ must belong to one maximal class in $\mathfrak{MC}$. Denote the maximal class containing $i$ as $\mathcal{MC}_{i}$. Then by Proposition 7,
$\mathcal{MC}_{i}\subseteq\mathcal{C}_{i}.$
Assume $k\in\mathcal{C}_{i}\backslash\mathcal{MC}_{i}$. Note that $|\mathcal{MC}_{i}|>1$ because the graph is weakly connected. Then $\exists l\neq i\in\mathcal{MC}_{i}$ s.t. $\sigma_{kl}=0$, because otherwise $k$ belongs to the same maximal class with every node in $\mathcal{MC}_{i}$, in particular with the source node, meaning that $k\in\mathcal{MC}_{i}$, which is a contradiction. However, $k,l\in\mathcal{C}_{i}$, and $\sigma_{kl}=0$ means that $\mathcal{C}_{i}\in\mathfrak{C}^{\prime}$, which is also a contradiction. Therefore,
$\mathcal{MC}_{i}=\mathcal{C}_{i},$
contradicting the fact that $\mathcal{C}_{i}\notin\mathfrak{MC}$.
In conclusion,
$\mathfrak{MC}=\mathfrak{C}\backslash\mathfrak{C}^{\prime}$.
With Theorem 2, we design an Algorithm 3 which returns the set of maximal classes using the support of $\Sigma$.
Algorithm 3 Algorithm for the reconstruction of maximal classes via the support of $\Sigma$
Input: $\mbox{supp}(\Sigma)$
   Output: the list of maximal classes of $G$: $\mathfrak{MC}$
[…]
2:for all nodes $i\in[n]$ do
3:  $\mathcal{C}_{i}\leftarrow\{j\mid j\in[n],\sigma_{ij}\neq 0\}$
4:  add $\mathcal{C}_{i}$ to $\mathfrak{MC}$
5:for all $k,l\in[n]\ s.t.\ k<l$ and $\sigma_{kl}==0$ do
6:  for all $\mathcal{C}_{i}\in\mathfrak{MC}$ do
7:    if $k,l\in\mathcal{C}_{i}$ then
8:     $\mathfrak{MC}\leftarrow\mathfrak{MC}\backslash\mathcal{C}_{i}$
9:for all $i\in[|\mathfrak{MC}|]$ do $\triangleright$ At this point, there might be repetitive elements in $\mathfrak{MC}$
10:  for all $i<j\leq|\mathfrak{MC}|$ do
11:    if $\mathfrak{MC}[i]==\mathfrak{MC}[j]$ then
12:     $\mathfrak{MC}\leftarrow\mathfrak{MC}\backslash\mathfrak{MC}[i]$ $\triangleright$ Keep only one of the repetitive elements
13:return $\mathfrak{MC}$
Proposition 7 and Theorem 2 imply that there is a bijection between the support of the covariance matrix $\Sigma$ and the set of maximal classes of the underlying graph, meaning that the set of maximal classes contains all the information of the underlying graph that we can reconstruct from the support of $\Sigma$. Additionally, by the definition of maximal classes, we are able to reconstruct a list of possible graphs admitting this set of maximal classes. Therefore, based on the support of $\Sigma$, a list of possible graphs admitting the corresponding set of maximal classes can be reconstructed. The following example is an illustration of how we could apply the results above.
Example 6.
Let $n=3$, assume the support of $\Sigma$ is
$\Sigma=\left[\begin{array}[]{ccc}\ast&\ast&0\\ \ast&\ast&\ast\\ 0&\ast&\ast\end{array}\right],$
where $\ast$ represents a non-zero element. In this case,
$\mathfrak{C}=\left\{\{1,2\},\{1,2,3\},\{2,3\}\right\},$
and
$\mathfrak{C}^{\prime}=\left\{\{1,2,3\}\right\},$
because $\sigma_{13}=0$. Hence, the set of maximal classes is
$\mathfrak{MC}=\mathfrak{C}\backslash\mathfrak{C}^{\prime}=\left\{\{1,2\},\{2,3\}\right\}.$
Note that because the node $2$ is in both of the two maximal classes, then it cannot be a source node in either of them, because otherwise there would have been only one maximal class with source $2$. Hence, $1$ and $3$ are the sources in respective maximal classes. Figure 3 shows the only possible graph $G=\left(V,\mathfrak{E}_{G}\right)$ associated with this $\Sigma$, i.e. $V=\{1,2,3\}$, $\mathfrak{E}_{G}=\left\{(1,1),(2,2),(3,3),(1,2),(3,2)\right\}$.
Figure 3: The graph associated with maximal classes $\left\{\{1,2\},\{2,3\}\right\}$.
5 Identifiability results
This section presents identifiability results for stationary VAR(1) models, which are also the main contributions of this paper. The results are divided into two parts: one of them is based on maximal classes (Section 5.1), and the other one examines the dimensions of the models (Section 5.2). Moreover, we provide a summary and illustration of the results at the end of this section.
5.1 Maximal class characterization
In this section, we propose two criteria for generic identifiability using maximal classes, one of which only applies to models with the same dimension, while the other one is generialized to models with any dimension. Before presenting the main results, some lemmas and propositions describing the relationship between maximal classes and the parameters of the model as well as the rank of the Jacobian matrix are introduced. These results are not only necessary in proving the main theorems, but also strengthen our understanding of the nature of maximal classes and the Jacobian matrix.
Lemma 6.
Let $\mathbf{M}_{G}$ be a stationary VAR(1) model with corresponding directed graph $G=\left(V,\mathfrak{E}_{G}\right)$, then for any two nodes $i,j\in[n]$, $\left(\Sigma\Lambda\right)_{ij}=\left(\Sigma\Lambda\right)_{ji}=0$ if and only if $i$ and $j$ do not belong to the same maximal class of $G$.
Proof.
By the proof of Proposition 7, we know that for all $a,b\in V$,
$\sigma_{ab}=\omega\sum_{k=0}^{+\infty}\sum_{l=1}^{n}\lambda_{la}^{\left[k\right]}\lambda_{lb}^{\left[k\right]}.$
Therefore, if $i$ and $j$ do not belong to the same maximal class,
$\left(\Sigma\Lambda\right)_{ij}=\sum_{s=1}^{n}\sigma_{is}\lambda_{sj}=\omega\sum_{s=1}^{n}\sum_{k=0}^{+\infty}\sum_{l=1}^{n}\lambda_{li}^{\left[k\right]}\lambda_{ls}^{\left[k\right]}\lambda_{sj}=0.$
Because if $\left(\Sigma\Lambda\right)_{ij}\neq 0$, then $\exists k\in\mathbb{N},s,l\in V$ s.t. $\lambda_{li}^{\left[k\right]}\lambda_{ls}^{\left[k\right]}\lambda_{sj}\neq 0$, i.e. there exist the following two paths in $G$:
$l\rightsquigarrow i,\mbox{ and }l\rightsquigarrow s\rightarrow j,$
where $\rightsquigarrow$ represents a directed path, not necessarily a single edge. In this case, $i$ and $j$ have the same ancestor $l$, thus belong to the same maximal class, which contradicts the assumption. Similarly,
$\left(\Sigma\Lambda\right)_{ji}=0.$
On the other hand, if $\left(\Sigma\Lambda\right)_{ij}=\omega\sum_{s=1}^{n}\sum_{k=0}^{+\infty}\sum_{l=1}^{n}\lambda_{li}^{\left[k\right]}\lambda_{ls}^{\left[k\right]}\lambda_{sj}=0$, then by Lemma 5, generically,
$\forall k\in\mathbb{N},s,l\in\left[n\right],\lambda_{li}^{\left[k\right]}\lambda_{ls}^{\left[k\right]}\lambda_{sj}=0,$
Let $s=j$, since $\lambda_{jj}\neq 0$,
$\forall k\in\mathbb{N},l\in\left[n\right],\lambda_{li}^{\left[k\right]}\lambda_{ls}^{\left[k\right]}=0.$
By Lemma 5, it implies that $i$ and $j$ do not have the same ancestor, and they are not connected by directed paths, i.e. $i$ and $j$ do not belong to the same maximal class. ∎
The following proposition highlights the link between maximal classes and the null columns of the Jacobian matrix, which is the foundation of characterizing models using maximal classes.
Proposition 8.
Let $\mathbf{M}_{G}$ be a stationary VAR(1) model with a corresponding directed graph $G=\left(V,\mathfrak{E}_{G}\right)$, then the Jacobian matrix satisfies: $\forall i\leq j\in[n]$, $\mathbf{J}_{G}^{\left[\sigma_{ij}\right]}=\mathbf{0}$ if and only if $i$ and $j$ do not belong to the same maximal class. Here $\mathbf{J}_{G}^{\left[\sigma_{ij}\right]}$ represents the column of $\mathbf{J}_{G}$ that corresponds to $\sigma_{ij}$.
Proof.
Firstly, we prove that if $i$ and $j$ do not belong to the same maximal class, then $\mathbf{J}_{G}^{\left[\sigma_{ij}\right]}=\mathbf{0}$. By Proposition 7 and Lemma 6, we know that
$\sigma_{ij}=\sigma_{ji}=0,\mbox{ and }\left(\Sigma\Lambda\right)_{ij}=\left(\Sigma\Lambda\right)_{ji}=0.$
Recall from Theorem 1 that
$\overline{\mathbf{J}}_{G}=\psi_{G}\left(B\right)\left(I_{n^{2}}-\Lambda\otimes\Lambda\right)^{-1},$
where
$B$ $=\left[\begin{array}[]{c}\left(\Sigma\Lambda\otimes I_{n}\right)\left(I_{n^{2}}+\mathbf{P}\right)\\ \mbox{vec}\left(I_{n^{2}}\right)^{T}\end{array}\right].$
Therefore,
$\mathbf{J}_{G}^{\left[\sigma_{ij}\right]}=\overline{\mathbf{J}}_{G}^{\left[\sigma_{ij}\right]}=\psi_{G}\left(B\right)\left(\left(I_{n^{2}}-\Lambda\otimes\Lambda\right)^{-1}\right)^{\left[\sigma_{ij}\right]},$
where
$\left(\left(I_{n^{2}}-\Lambda\otimes\Lambda\right)^{-1}\right)^{\left[\sigma_{ij}\right]}=\sum_{k=0}^{+\infty}\left[\begin{array}[]{cccccccccc}\lambda_{1i}^{\left[k\right]}\lambda_{1j}^{\left[k\right]}&\cdots&\lambda_{1i}^{\left[k\right]}\lambda_{nj}^{\left[k\right]}&\lambda_{2i}^{\left[k\right]}\lambda_{1j}^{\left[k\right]}&\cdots&\lambda_{2i}^{\left[k\right]}\lambda_{nj}^{\left[k\right]}&\cdots&\lambda_{ni}^{\left[k\right]}\lambda_{1j}^{\left[k\right]}&\cdots&\lambda_{ni}^{\left[k\right]}\lambda_{nj}^{\left[k\right]}\end{array}\right]^{T}.$
Recall that $\psi_{G}$ removes the rows in $\overline{\mathbf{J}}$ that correspond to $\lambda_{ab}$, where $\left(a,b\right)\notin\mathfrak{E}_{G}$. For all $s,t\in\left[n\right]$ s.t. $\left(s,t\right)\in\mathfrak{E}_{G}$, denote the row in $\psi_{G}\left(B\right)$ that corresponds to $\lambda_{st}$ as $\psi_{G}\left(B\right)^{\left(\lambda_{st}\right)}$, and the element in $\mathbf{J}_{G}^{\left[\sigma_{ij}\right]}$ that corresponds to $\lambda_{st}$ as $\mathbf{J}_{G}^{\left[\lambda_{st},\sigma_{ij}\right]}$. Then
$\psi_{G}\left(B\right)^{\left(\lambda_{st}\right)}=$ $\left[\begin{array}[]{cccc}\begin{array}[]{ccccc}0&\cdots&\left(\Sigma\Lambda\right)_{s1}&\cdots&0\end{array}&\cdots&\begin{array}[]{ccccc}\left(\Sigma\Lambda\right)_{s1}&\cdots&\left(\Sigma\Lambda\right)_{st}&\cdots&\left(\Sigma\Lambda\right)_{sn}\end{array}$$\cdots\end{array}\right$.
$\left.\begin{array}[]{ccccc}0&\cdots&\left(\Sigma\Lambda\right)_{sn}&\cdots&0\end{array}\right].$
Therefore,
$\mathbf{J}_{G}^{\left[\lambda_{st},\sigma_{ij}\right]}$ $=\psi_{G}\left(B\right)^{\left(\lambda_{st}\right)}\left(\left(I_{n^{2}}-\Lambda\otimes\Lambda\right)^{-1}\right)^{\left[\sigma_{ij}\right]}$
$=\sum_{k=0}^{+\infty}\left(\sum_{a=1}^{n}\left(\Sigma\Lambda\right)_{sa}\lambda_{ai}^{\left[k\right]}\lambda_{tj}^{\left[k\right]}+\sum_{b=1}^{n}\left(\Sigma\Lambda\right)_{sb}\lambda_{ti}^{\left[k\right]}\lambda_{bj}^{\left[k\right]}\right).$
Consider the first term of the sum: for all $a\in\left[n\right]$ and $k\in\mathbb{N}$, if $\left(\Sigma\Lambda\right)_{sa}\lambda_{ai}^{\left[k\right]}\lambda_{tj}^{\left[k\right]}\neq 0$, then from the fact that $\left(\Sigma\Lambda\right)_{sa}\neq 0$, we know by Lemma 6 that $s$ and $a$ belong to the same maximal class, i.e. there is a directed path between $s$ and $a$ (Case 1), or there exists a node $l\neq s,a\in\left[n\right]$ s.t. there are directed paths from $l$ to $s$ and $a$ respectively (Case 2). Additionally, there exist direct paths among the following nodes:
$s\rightarrow t\rightsquigarrow j\mbox{, and }a\rightsquigarrow i.$
  • Case 1.
    without loss of generality, assume the path is from $s$ to $a$, then there exist directed paths: $s\rightsquigarrow t\rightsquigarrow j$, and $s\rightsquigarrow a\rightsquigarrow i$. It contradicts the assumption.
  • Case 2.
    assume there exist directed paths: $l\rightsquigarrow s$, and $l\rightsquigarrow a$, then there exist directed paths: $l\rightsquigarrow a\rightsquigarrow i$, and $l\rightsquigarrow s\rightarrow t\rightsquigarrow j$. It also contradicts the assumption.
Therefore, for all $a\in\left[n\right]$ and $k\in\mathbb{N}$, $\left(\Sigma\Lambda\right)_{sa}\lambda_{ai}^{\left[k\right]}\lambda_{tj}^{\left[k\right]}=0$. The second term of the sum is also 0 by similar arguments. Hence, for all $s,t\in\left[n\right]$ s.t. $\left(s,t\right)\in\mathfrak{E}_{G}$,
$\mathbf{J}_{G}^{\left[\lambda_{st},\sigma_{ij}\right]}=0.$
In addition,
$\mathbf{J}_{G}^{\left[\omega,\sigma_{ij}\right]}=\mbox{vec}\left(I_{n^{2}}\right)\left(\left(I_{n^{2}}-\Lambda\otimes\Lambda\right)^{-1}\right)^{\left[\sigma_{ij}\right]}=\sum_{k=0}^{+\infty}\sum_{a=1}^{n}\lambda_{ai}^{\left[k\right]}\lambda_{aj}^{\left[k\right]}=0,$
because if there exists $k\in\mathbb{N}$ and $a\in\left[n\right]$ s.t. $\lambda_{ai}^{\left[k\right]}\lambda_{aj}^{\left[k\right]}\neq 0$, then there exist directed paths: $a\rightsquigarrow i$ and $a\rightsquigarrow j$, which contradicts the assumption.
In conclusion,
$\mathbf{J}_{G}^{\left[\sigma_{ij}\right]}=\mathbf{0}.$
Next, we prove that if $i$ and $j$ belong to the same maximal class, i.e. there is a directed path between $i$ and $j$, or there exists $k\neq i,j\in\left[n\right]$ s.t. there are directed paths from $k$ to $i$ and $j$ respectively, then $\mathbf{J}_{G}^{\left[\sigma_{ij}\right]}\neq\mathbf{0}$.
In particular,
$\mathbf{J}_{G}^{\left[\omega,\sigma_{ij}\right]}=\mbox{vec}\left(I_{n^{2}}\right)\left(\left(I_{n^{2}}-\Lambda\otimes\Lambda\right)^{-1}\right)^{\left[\sigma_{ij}\right]}=\sum_{k=0}^{+\infty}\sum_{a=1}^{n}\lambda_{ai}^{\left[k\right]}\lambda_{aj}^{\left[k\right]}\neq 0,$
because if $\mathbf{J}_{G}^{\left[\omega,\sigma_{ij}\right]}=0$, then by Lemma 5, for all $k\in\left[n\right]$ and $a\in\left[n\right]$, $\lambda_{ai}^{\left[k\right]}\lambda_{aj}^{\left[k\right]}=0$. It means that there is no directed path from $a$ to $i$ and $j$ respectively (here $a$ may equal to $i$ or $j$), which contradicts the assumption. Therefore, $\mathbf{J}_{G}^{\left[\omega,\sigma_{ij}\right]}\neq 0$, and moreover,
$\mathbf{J}_{G}^{\left[\sigma_{ij}\right]}\neq\mathbf{0}.$
Proposition 8 shows that two nodes that do not belong to the same maximal class will result in a null column in the Jacobian matrix. This is already an important piece of information on the Jacobian matroid. With one more proposition on the rank of the Jacobian matrix, we are able to prove the first two main results of this paper.
Lemma 7.
Let $\mathbf{M}_{G}$ be a stationary VAR(1) model with an associated directed graph $G$. Then
$rank\left(\mathbf{J}_{G}\right)=rank\left(\psi_{G}\left(B\right)\right).$
Proof.
The proof is given in Appendix H. ∎
Lemma 7 ensures that when we need to study the rank of $\mathbf{J}_{G}$, we can look at the simpler matrix $\psi_{G}(B)$ instead. The following proposition provides an upper and lower bound for the rank of the Jacobian matrix based on this property.
Proposition 9.
Let $\mathbf{M}_{G}$ be a stationary VAR(1) model. Then
$n\leq rank\left(\mathbf{J}_{G}\right)\leq\min\left\{n_{r},n_{c}^{\prime}\right\},$
where
$n_{r}$ $=E_{G}+1;$
$n_{c}^{\prime}$ $=|\left\{\{a,b\}\mid a,b\in\left[n\right],a,b\mbox{ belong to the same maximal class}\right\}|.$
Proof.
Consider a $n\times n$ submatrix of $\psi_{G}(B)$, with rows and columns corresponding to $\lambda_{ii}$ and $\sigma_{jj}$ respectively, for all $i,j\in[n]$. Then it is a diagonal matrix with non-zero entries on the diagonal. Therefore, the rank of this submatrix is $n$, and the rank of $\psi_{G}(B)$ is at least $n$ by the Guttman rank additivity formula ([8]). The result follows from the fact that $\mbox{rank}\left(\mathbf{J}_{G}\right)=\mbox{rank}\left(\psi_{G}(B)\right)$.
By definition, $\mathbf{J}_{G}$ is of size $\left(E_{G}+1\right)\times\frac{n\left(n+1\right)}{2}$. So
$rank\left(\mathbf{J}_{G}\right)\leq\min\left\{n_{r},\frac{n\left(n+1\right)}{2}\right\}.$
In fact, some of the columns of $\mathbf{J}_{G}$ are zero vectors. We only count the non-zero columns, which is $n_{c}^{\prime}$ by Proposition 8. Therefore,
$rank\left(\mathbf{J}_{G}\right)\leq\min\left\{n_{r},n_{c}^{\prime}\right\}.$
Now, we present the first main result that characterizes models with the same dimension using maximal classes, based on the lemmas and propositions introduced above.
Theorem 3.
Let $\left\{\mathbf{M}_{k}\right\}_{k=1}^{K}$ be a finite set of stationary VAR(1) models corresponding to graphs $\left\{G_{k}\right\}_{k=1}^{K}$. If the models have the same dimension and for all distinct pairs $\left(k_{1},k_{2}\right)$ of values from $1$ to $K$, $G_{k_{1}}$ and $G_{k_{2}}$ have different maximal classes, then these models (or the discrete parameter $k$) are generically identifiable.
Proof.
Let $\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ be two stationary VAR(1) models with the same dimension and different maximal classes. Then there exist two nodes $a,b\in[n]$ s.t. $a$ and $b$ belong to the same maximal class in $G_{1}$ and do not belong to the same maximal class in $G_{2}$, because otherwise by Proposition 7 and Theorem 2, they have the same set of maximal classes. Therefore, by Proposition 8,
$\mathbf{J}_{G_{1}}^{\left[\sigma_{ab}\right]}\neq\mathbf{0}\mbox{, and }\mathbf{J}_{G_{2}}^{\left[\sigma_{ab}\right]}=\mathbf{0}.$
Let $A,B\in\mathbb{R}^{n}$. The notation $A\perp\!\!\!\perp B$ indicates that $A$ and $B$ are linearly independent, and $A\not\!\perp\!\!\!\perp B$ indicates that $A$ and $B$ are not linearly independent. Then, there exist $a^{\prime},b^{\prime}\in[n]$ where $\{a^{\prime},b^{\prime}\}\neq\{a,b\}$ s.t.
$\mathbf{J}_{G_{1}}^{\left[\sigma_{ab}\right]}\perp\!\!\!\perp\mathbf{J}_{G_{1}}^{\left[\sigma_{a^{\prime}b^{\prime}}\right]},$
because otherwise for all $a^{\prime},b^{\prime}\in[n]$ where $\{a^{\prime},b^{\prime}\}\neq\{a,b\}$, we have
$\mathbf{J}_{G_{1}}^{\left[\sigma_{ab}\right]}\not\!\perp\!\!\!\perp\mathbf{J}_{G_{1}}^{\left[\sigma_{a^{\prime}b^{\prime}}\right]},$
meaning that $\mbox{rank}\left(\mathbf{J}_{G_{1}}\right)\leq 1$, which contradicts the fact that $\mbox{rank}\left(\mathbf{J}_{G_{1}}\right)\geq n\geq 2$ (see Proposition 9). Additionally, since $\mathbf{J}_{G_{2}}^{\left[\sigma_{ab}\right]}=\mathbf{0}$, we have
$\mathbf{J}_{G_{2}}^{\left[\sigma_{ab}\right]}\not\!\perp\!\!\!\perp\mathbf{J}_{G_{2}}^{\left[\sigma_{a^{\prime}b^{\prime}}\right]}.$
Combining the results above, we have found a pair of columns of the Jacobian matrix corresponding to $\sigma_{ab}$ and $\sigma_{a^{\prime}b^{\prime}}$ such that they are linearly independent in $\mathbf{M}_{1}$, and not in $\mathbf{M}_{2}$. Therefore, $\mathcal{J}\left(\mathbf{M}_{1}\right)\neq\mathcal{J}\left(\mathbf{M}_{2}\right)$, and hence by Proposition 3, the two models are generically identifiable. ∎
Note that Theorem 3 is restricted to models with the same dimension, i.e. for models with possibly different dimensions, the theorem does not hold. The following theorem states that this constraint could be released by having more constraints on the maximal classes of the graphs.
Theorem 4.
Let $\left\{\mathbf{M}_{k}\right\}_{k=1}^{K}$ be a finite set of stationary VAR(1) models corresponding to graphs $\left\{G_{k}\right\}_{k=1}^{K}$. If for all distinct pairs $\left(k_{1},k_{2}\right)$ of values from $1$ to $K$, $G_{k_{1}}$ and $G_{k_{2}}$ satisfies the following two conditions:
1. There exist $i,j\in[n]$ s.t. $i,j$ belong to the same maximal class in $G_{k_{1}}$, but do not in $G_{k_{2}}$.
2. There exist $\in[n] s.t. s,t$s,t belong to the same maximal class in $G_{k_{2}}$, but do not in $G_{k_{1}}$.
then the models (or the discrete parameter $k$) are generically identifiable.
Proof.
Let $\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ be two stationary VAR(1) models corresponding to graphs $G_{1}$ and $G_{2}$ that satisfy the two conditions in the theorem. Then, from the proof of Theorem 3, we know that:
$\exists S_{1}\in\mathcal{J}\left(\mathbf{M}_{1}\right)\backslash\mathcal{J}\left(\mathbf{M}_{2}\right)\mbox{ and }\exists S_{2}\in\mathcal{J}\left(\mathbf{M}_{2}\right)\backslash\mathcal{J}\left(\mathbf{M}_{1}\right).$
Therefore, by Proposition 2, $\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ are generically identifiable, no matter the dimensions. ∎
Example 7.
Figure 4 presents two graphs that are generically identifiable using the results introduced in this section.
(a) $G_{1}$
(b) $G_{2}$
Figure 4: Identifiable graphs using Theorem 4
The sets of maximal classes $\mathfrak{MC}_{1}$ of $G_{1}$ and $\mathfrak{MC}_{2}$ of $G_{2}$ are:
$\mathfrak{MC}_{1}=\left\{\{1,2,3\},\{3,4\}\right\},\mbox{ and }\mathfrak{MC}_{2}=\left\{\{1,2\},\{2,3,4\}\right\}.$
Because at this point, we don’t know the dimension of the models, we can only apply Theorem 4. In fact, nodes $1$ and $3$ belong to the same maximal class in $G_{1}$, but not in $G_{2}$. Similarly, nodes $2$ and $4$ belong to the same maximal class in $G_{2}$, but not in $G_{1}$. By Theorem 4, the two graphs are generically identifiable.
While Theorems 3 and 4 already provide criteria for generic identifiability that are easy to check, if the dimension of the model could be calculated from the graph, we would only need to focus on the identifiability of models with the same dimension, and always apply the weaker condition on maximal classes introduced in Theorem 3. Indeed, the following section presents that if we consider only a subset of models whose associated graphs do not have ”multi-edges”, meaning that $(i,j)\in\mathfrak{E}_{G}$ implies that $(j,i)\notin\mathfrak{E}_{G}$, we are able to derive the exact dimension of the model from the graph.
5.2 The dimension of the model for graphs without multi-edges
This section demonstrates that for a subset of models whose corresponding graphs do not have ”multi-edges”, the dimension (i.e. the rank of the Jacobian matrix) can be determined from the graph. Using this result, we provide additional identifiability criteria based on the dimensions of the model, complementing those in Section 5.1. By integrating these results, we are able to expand the scope of models that are identifiable. On the other hand, for the cases where multi-edges often exist, such as in ecological research, the results in this section do not apply anymore.
The following assumption excludes ”multi-edges”, which is needed in the study of the rank of the Jacobian matrix. In particular, it ensures the existence of ”triplets” in Lemma 10, which is crucial in proving that the Jacobian matrix is of full rank in Theorem 5.
Assumption 1.
For all graphs $G$, we exclude the existence of ”multi-edges”, i.e. $\forall i\neq j\in[n]$, $(i,j)\in\mathfrak{E}_{G}$ implies $(j,i)\notin\mathfrak{E}_{G}$.
Theorem 5.
Let $\mathbf{M}_{G}$ be a stationary VAR(1) model that satisfies Assumption 1, then
$\mbox{rank}(\mathbf{J}_{G})=\min\left\{n_{r},n_{c}^{\prime}\right\},$ (5)
where $n_{r}$ and $n_{c}^{\prime}$ are defined in Proposition 9, is generically true, i.e. the set of values such that (5) is not true has Lebesgue measure zero.
Note that by definition, $n_{r}$ is the number of rows, and $n_{c}^{\prime}$ is the number of non-zero columns of $\mathbf{J}_{G}$. Thus Theorem 5 in fact indicates that all of $\mathbf{J}_{G}$’s non-zero columns form a full rank matrix.
Example 8.
Consider a stationary VAR(1) model with the corresponding graph below. The set of maximal classes $\mathfrak{MC}$, $n_{r}$ and $n_{c}^{\prime}$ are:
$\mathfrak{MC}=$ $\left\{\{1,2,3,4\},\{3,5\}\right\};$
$n_{r}=$ $E_{G}+1=5+4+1=10;$
$n_{c}^{\prime}=$ $|\left\{\{1,1\},\{2,2\},\{3,3\},\{4,4\},\{5,5\},\right$.
$\left.\{1,2\},\{1,3\},\{1,4\},\{2,3\},\{2,4\},\{3,4\},\{3,5\}\right\}|=12.$
Therefore,
${\operatorname{dim}}\left(\mathbf{M}_{G}\right)=\mbox{rank}\left(\mathbf{J}_{G}\right)=\min\left\{n_{r},n_{c}^{\prime}\right\}=10.$
The proof of Theorem 5 is technical, and aligns with generic settings throughout this paper. It is divided into two separate cases: $n_{r}\leq n_{c}^{\prime}$ and $n_{r}>n_{c}^{\prime}$. We only introduce the proof for the first case, since the proof of the second case resembles the first one. For a complete proof of the second case, see Appendix J.
Now, assume $n_{r}\leq n_{c}^{\prime}$, and the goal is to prove that under this condition, $rank\left(\mathbf{J}_{G}\right)=n_{r}$. For this, Lemmas 8-11 are needed. They transform the rank of the Jacobian matrix to that of another matrix that is derived from the graph, and prove that this matrix is full rank generically, based on a special graphical structure that exists in all graphs in this case.
Let $\mathfrak{E}_{G}^{\prime}$ be the set of edges excluding self-loops, and $E_{G}^{\prime}=|\mathfrak{E}_{G}^{\prime}|$. Define the set of pairs of nodes such that they are not directly connected by one edge, but belong to the same maximal class:
[…]
and denote the cardinality of $\mathfrak{C}_{G}^{mc}$ as $C_{G}^{mc}$. Under Assumption 1, the following lemma introduces a necessary and sufficient condition for $n_{r}\leq n_{c}^{\prime}$.
Lemma 8.
Let $\mathbf{M}_{G}$ be a stationary VAR(1) model that satisfies Assumption 1, then $n_{r}\leq n_{c}^{\prime}$ if and only if there exist two nodes $k,l\in\left[n\right]$, s.t. $\{k,l\}\in\mathfrak{C}_{G}^{mc}$.
Proof.
Assume $\forall k,l\in[n]$, $\{k,l\}\notin\mathfrak{C}_{G}^{mc}$. Then either $(k,l)$ or $(l,k)\in\mathfrak{E}_{G}$ or $k,l$ do not belong to the same maximal class, i.e. $(\Sigma\Lambda)_{kl}=(\Sigma\Lambda)_{lk}=0$. In this case, $n_{c}^{\prime}=E_{G}$ because for all pairs of nodes $i,j\in[n]$, $(\Sigma\Lambda)_{ij}\neq 0$ if and only if $\left(i,j\right)$ or $\left(j,i\right)\in\mathfrak{E}_{G}$ (but never at the same time because of Assumption 1). Since $n_{r}=E_{G}+1$, $n_{r}>n_{c}^{\prime}$, which contradicts the premise. ∎
Lemma 8 indicates that under Assumption 1, the graphs that don’t satisfy $n_{r}\leq n_{c}^{\prime}$ are the ones such that there are at least two maximal classes, and the undirected subgraph of each maximal class is complete.
Next, we introduce a new matrix $B_{G}$ derived from the Jacobian matrix. And it will be shown in Lemma 9 that the Jacobian matrix is full rank if the matrix $B_{G}$ is full rank.
Definition 10.
Let $\mathbf{M}_{G}$ be a stationary VAR(1) model that satisfies the condition $n_{r}\leq n_{c}^{\prime}$. Then consider a square matrix of size $E_{G}^{\prime}+1$, where the rows correspond to $\lambda_{ij}$, s.t. $\left(i,j\right)\in\mathfrak{E}_{G}^{\prime}$, and $\omega$, and the columns correspond to $\sigma_{S}$, where
$S=\left\{\{i,j\}\mid(i,j)\in\mathfrak{E}_{G}^{\prime}\right\}\cup\{\{k,l\}\},\mbox{ for any }\{k,l\}\in\mathfrak{C}_{G}^{mc}.$
This matrix, denoted as $B_{G}$, is defined as
$B_{G}^{\left[\lambda_{ij},\sigma_{ab}\right]}$ $=\delta_{ja}\left[\left(\Sigma\Lambda\right)_{ib}\left(\Sigma\Lambda\right)_{jj}-\left(\Sigma\Lambda\right)_{ij}\left(\Sigma\Lambda\right)_{jb}\right]+\delta_{jb}\left[\left(\Sigma\Lambda\right)_{ia}\left(\Sigma\Lambda\right)_{jj}-\left(\Sigma\Lambda\right)_{ij}\left(\Sigma\Lambda\right)_{ja}\right];$
$B_{G}^{\left[\omega,\sigma_{ab}\right]}$ $=\left(\Sigma\Lambda\right)_{ab}\left(\Sigma\Lambda\right)_{aa}^{-1}+\left(\Sigma\Lambda\right)_{ba}\left(\Sigma\Lambda\right)_{bb}^{-1},$
where $a,b\in[n]$ and $\{a,b\}\in S$.
Note that given a directed graph $G$, $B_{G}$ might not be unique, because when $C_{G}^{mc}>1$, there are multiple options for $S$.
Example 9.
Consider a stationary VAR(1) model $\mathbf{M}_{G}$ with a corresponding directed graph $G=\left(V,\mathfrak{E}_{G}\right)$, where $V=\left\{1,2,3\right\}$ and $\mathfrak{E}_{G}=\left\{\left(1,2\right),\left(2,3\right)\right\}$. Here there exists only one $B_{G}$, which is
$B_{G}=\left[\begin{array}[]{ccc}\left(\Sigma\Lambda\right)_{11}\left(\Sigma\Lambda\right)_{22}-\left(\Sigma\Lambda\right)_{12}\left(\Sigma\Lambda\right)_{21}&\left(\Sigma\Lambda\right)_{13}\left(\Sigma\Lambda\right)_{22}-\left(\Sigma\Lambda\right)_{12}\left(\Sigma\Lambda\right)_{23}&0\\ 0&\left(\Sigma\Lambda\right)_{22}\left(\Sigma\Lambda\right)_{33}-\left(\Sigma\Lambda\right)_{23}\left(\Sigma\Lambda\right)_{32}&\left(\Sigma\Lambda\right)_{21}\left(\Sigma\Lambda\right)_{33}-\left(\Sigma\Lambda\right)_{23}\left(\Sigma\Lambda\right)_{31}\\ \left(\Sigma\Lambda\right)_{12}\left(\Sigma\Lambda\right)_{11}^{-1}+\left(\Sigma\Lambda\right)_{21}\left(\Sigma\Lambda\right)_{22}^{-1}&\left(\Sigma\Lambda\right)_{23}\left(\Sigma\Lambda\right)_{22}^{-1}+\left(\Sigma\Lambda\right)_{32}\left(\Sigma\Lambda\right)_{33}^{-1}&\left(\Sigma\Lambda\right)_{13}\left(\Sigma\Lambda\right)_{11}^{-1}+\left(\Sigma\Lambda\right)_{31}\left(\Sigma\Lambda\right)_{33}^{-1}\end{array}\right].$
Lemma 9.
Let $\mathbf{M}_{G}$ be a stationary VAR(1) model that satisfies Assumption 1 and the condition $n_{r}\leq n_{c}^{\prime}$. Then $\mbox{rank}\left(\mathbf{J}_{G}\right)=n_{r}$ if there exists a $B_{G}$ that is full rank, i.e. $\mbox{rank}\left(B_{G}\right)=E_{G}^{\prime}+1$.
Sketch of proof.
By Lemma 7, $\mbox{rank}\left(\mathbf{J}_{G}\right)=\mbox{rank}\left(\psi_{G}(B)\right)$, and therefore it is sufficient to prove that $\mbox{rank}\left(\psi_{G}(B)\right)=n_{r}$. Consider a $\left(E_{G}+1\right)\times\left(E_{G}+1\right)$ submatrix of $\psi_{G}\left(B\right)$, where the columns correspond to the set
$\left\{\sigma_{ij}\mid i,j\in\left[n\right],(i,j)\mbox{ or }(j,i)\in\mathfrak{E}_{G}\right\}\cup\left\{\sigma_{kl}\right\},$
where $\left\{k,l\right\}\in\mathfrak{C}_{G}^{mc}$. Note that $\mathfrak{C}_{G}^{mc}\neq\emptyset$ by Lemma 8. In fact, this submatrix is a block matrix after reordering rows and columns. Apply the the Guttman rank additivity formula (see [8]), this matrix is full rank if and only if $B_{G}$ is full rank, up to a multiplication of a diagonal matrix with non-zero diagonals. For complete proof, see Appendix I. ∎
Next, we prove that under Assumption 1 and the condition $n_{r}\leq n_{c}^{\prime}$, $B_{G}$ is generically full rank, i.e. the Jacobian matrix $\mathbf{J}_{G}$ is full rank.
Lemma 10.
Let $\mathbf{M}_{G}$ be a stationary VAR(1) model that satisfies the condition $n_{r}\leq n_{c}^{\prime}$ and Assumption 1. Then there exists three nodes: $k,a,l\in\left[n\right]$ s.t. one of the following circumstances holds:
1. $\left(k,a\right),\left(a,l\right)\in\mathfrak{E}_{G}^{\prime}$ and $\left(k,l\right),\left(l,k\right)\notin\mathfrak{E}_{G}^{\prime}$.
2. $\left(a,k\right),\left(a,l\right)\in\mathfrak{E}_{G}^{\prime}$ and $\left(k,l\right),\left(l,k\right)\notin\mathfrak{E}_{G}^{\prime}$.
Proof.
From Lemma 8, we know that there exist two nodes $k,l\in[n]$ such that $\{k,l\}\in\mathfrak{C}_{G}^{mc}$, i.e. $(k,l),(l,k)\notin\mathfrak{E}_{G}^{\prime}$, but $k$ and $l$ belong to the same maximal class. This ensures the existence of the triplets. ∎
The following lemma states that the set of values such that $B_{G}$ is not full rank has zero Lebesgue measure. We keep the proof in the main text because it again aligns with the generic settings of this paper, and uses the existence of the ”triplet” structure introduced in Lemma 10, which is technical and insightful.
Lemma 11.
Let $\mathbf{M}_{G}$ be a stationary VAR(1) model that satisfies Assumption 1 and $n_{r}\leq n_{c}^{\prime}$. Define a subset $M_{G}^{B}$ of $M_{G}$:
$M_{G}^{B}:=\left\{\Lambda\in M_{G}\mid\det\left(B_{G}\right)=0\right\},$
then
$\mu_{G}\left(M_{G}^{B}\right)=0,$
where $\mu_{G}$ is the same measure defined in Proposition 4.
Proof.
First, we prove that the function
$f_{G}:M_{G}\times\mathbb{R}^{+}$ $\rightarrow\mathbb{R}$
$\left(\Lambda,\omega\right)$ $\mapsto\det\left(B_{G}\right)$
is rational. From Appendix B, we know that the parametrization map
$\phi_{G}:M_{G}\times\mathbb{R}^{+}$ $\rightarrow M_{n}\left(\mathbb{R}\right)$
$\left(\Lambda,\omega\right)$ $\mapsto\Sigma,\mbox{ s.t. }\mbox{vec}\left(\Sigma\right)=\left(I_{n}-\Lambda^{T}\otimes\Lambda^{T}\right)^{-1}\mbox{vec}\left(\omega I_{n}\right),$
is rational, i.e. $\Sigma$ is a matrix whose elements are Quotients of Polynomials (QOPs) with respect to the elements of $\Lambda$ and $\omega$. Therefore, $\left(\Sigma\Lambda\right)$ is also a matrix whose elements are QOPs. By definition, the elements of $B_{G}$ are the elements of $\left(\Sigma\Lambda\right)$ after summations, multiplications, and inversions. Hence, they are also QOPs. Since the determinant is also a result of summations and multiplications of the elements of the matrix, $f_{G}$ is rational, i.e. $\exists g,h\in\mathbb{R}\left[\left(\lambda_{ij}\mid\left(i,j\right)\in\mathfrak{E}_{G}\right),\omega\right]$ polynomials s.t.
$f_{G}=\frac{g}{h}.$
Next, we prove that the set of parameters such that $h=0$ has Lebesgue measure zero. Recall from (3), for all $i,j\in\left[n\right]$,
$\left(\Sigma\Lambda\right)_{ij}$ $=\sum_{k=1}^{n}\sigma_{ik}\lambda_{kj}=\sum_{k=1}^{n}\left(\mbox{vec}\left(\Sigma\right)\right)_{\left(k-1\right)n+i}\lambda_{kj}$
$=\omega\sum_{k=1}^{n}\sum_{l=1}^{n}\left(I_{n^{2}}-\Lambda^{T}\otimes\Lambda^{T}\right)^{-1}_{\left(k-1\right)n+i,\left(l-1\right)n+l}\lambda_{kj}.$
Since
$\left(I_{n^{2}}-\Lambda^{T}\otimes\Lambda^{T}\right)^{-1}=\frac{1}{\det\left(I_{n^{2}}-\Lambda^{T}\otimes\Lambda^{T}\right)}\mbox{adj}\left(I_{n^{2}}-\Lambda^{T}\otimes\Lambda^{T}\right),$
$\left(\Sigma\Lambda\right)_{ij}=\frac{P}{Q},$
where $\mbox{adj}(\cdot)$ is the adjugate matrix, $P\in\mathbb{R}\left[\left(\lambda_{ij}\mid\left(i,j\right)\in\mathfrak{E}_{G}\right),\omega\right]$, and $Q=\det\left(I_{n^{2}}-\Lambda^{T}\otimes\Lambda^{T}\right)\in\mathbb{R}\left[\lambda_{ij}\mid\left(i,j\right)\in\mathfrak{E}_{G}\right]$. Therefore $\exists m\in\mathbb{N}$ s.t.
$h=Q^{m}\prod_{i=1}^{n}\left(\Sigma\Lambda\right)_{ii}.$
From the fact that $\left(\Sigma\Lambda\right)_{ii}\neq 0$ for all $i\in\left[n\right]$,
$h=0\Leftrightarrow Q=0\Leftrightarrow\det\left(I_{n^{2}}-\Lambda^{T}\otimes\Lambda^{T}\right)=0\Leftrightarrow\det\left(I_{n^{2}}-\Lambda\otimes\Lambda\right)=0.$
Therefore by Proposition 4,
$\mu_{G}\left(\left\{\Lambda\in M_{G}\mid h=0\right\}\right)=\mu_{G}\left(M_{G}^{0}\right)=0.$
Finally, we prove that the set of parameters such that $g=0$ has Lebesgue measure zero. Since $\det\left(B_{G}\right)$ is a QOP, it’s sufficient to prove that for any directed graph $G=\left(V,\mathfrak{E}_{G}\right)$ that satisfies the premises, there exists $\Lambda_{0}\in M_{G}$, s.t. $\det\left(B_{G_{0}}\right)\neq 0$, where $B_{G_{0}}$ is $B_{G}$ with $\Lambda=\Lambda_{0}$. From Lemma 10, there exists three nodes $a,k,l\in\left[n\right]$, s.t. either $\left(k,a\right),\left(a,l\right)\in\mathfrak{E}_{G}^{\prime}$ and $\left(k,l\right),\left(l,k\right)\notin\mathfrak{E}_{G}^{\prime}$ (Case 1), or $\left(a,k\right),\left(a,l\right)\in\mathfrak{E}_{G}^{\prime}$ and $\left(k,l\right),\left(l,k\right)\notin\mathfrak{E}_{G}^{\prime}$ (Case 2). In both scenarios, $k,l$ belong to the same maximal class. Consider $B_{G}$ with rows correspond to $\lambda_{ij}$, where $\left(i,j\right)\in\mathfrak{E}_{G}^{\prime}$ and $\omega$, and columns correspond to $\sigma_{ab}$, where $\left(a,b\right)\in\mathfrak{E}_{G}^{\prime}$, and $\sigma_{kl}$. We discuss the two cases separately.
Case 1.
Consider $\Lambda_{0}=\left(\lambda_{ij}^{0}\right)$ that satisfies:
$\lambda_{ij}^{0}=\left\{\begin{array}[]{lcl}\lambda_{ij}^{0}\neq 0&,&i=j\mbox{ or }\left(i,j\right)=\left(k,a\right)\mbox{ or }\left(i,j\right)=\left(a,l\right)\\ 0&,&\mbox{otherwise}\end{array}\right$.,
then
$\left(\Sigma_{0}\Lambda_{0}\right)_{ij}\neq 0\Leftrightarrow i=j\mbox{ or }\left\{i,j\right\}\in\left\{\left\{k,a\right\},\left\{a,l\right\},\left\{k,l\right\}\right\}.$
Order the rows and columns of $B_{G}$ such that the rows correspond to $\lambda_{ka}$, $\lambda_{al}$, $\omega$ and $\lambda_{ij}$, and the columns correspond to $\sigma_{ka}$, $\sigma_{al}$, $\sigma_{kl}$ and $\sigma_{ij}$ respectively, where $\left(i,j\right)$ represents all the edges in $\mathfrak{E}_{G}^{\prime}$ other than $\left(k,a\right)$ and $\left(a,l\right)$. Note that the columns are ordered along with the corresponding rows. Consider the matrix $B_{G_{0}}$, which is defined as $B_{G}$ with $\Lambda=\Lambda_{0}$, and define it as a block matrix:
$B_{G_{0}}=\left[\begin{array}[]{cc}A&B\\ C&D\end{array}\right],$
where $A$ is a $3\times 3$ matrix consisting of the first three rows and columns of $B_{G_{0}}$. Then
$A=$
$\left[\begin{array}[]{ccc}\left(\Sigma_{0}\Lambda_{0}\right)_{kk}\left(\Sigma_{0}\Lambda_{0}\right)_{aa}-\left(\Sigma_{0}\Lambda_{0}\right)_{ka}\left(\Sigma_{0}\Lambda_{0}\right)_{ak}&\left(\Sigma_{0}\Lambda_{0}\right)_{kl}\left(\Sigma_{0}\Lambda_{0}\right)_{aa}-\left(\Sigma_{0}\Lambda_{0}\right)_{ka}\left(\Sigma_{0}\Lambda_{0}\right)_{al}&0\\ 0&\left(\Sigma_{0}\Lambda_{0}\right)_{aa}\left(\Sigma_{0}\Lambda_{0}\right)_{ll}-\left(\Sigma_{0}\Lambda_{0}\right)_{al}\left(\Sigma_{0}\Lambda_{0}\right)_{la}&\left(\Sigma_{0}\Lambda_{0}\right)_{ak}\left(\Sigma_{0}\Lambda_{0}\right)_{ll}-\left(\Sigma_{0}\Lambda_{0}\right)_{al}\left(\Sigma_{0}\Lambda_{0}\right)_{lk}\\ \left(\Sigma_{0}\Lambda_{0}\right)_{ka}\left(\Sigma_{0}\Lambda_{0}\right)_{kk}^{-1}+\left(\Sigma_{0}\Lambda_{0}\right)_{ak}\left(\Sigma_{0}\Lambda_{0}\right)_{aa}^{-1}&\left(\Sigma_{0}\Lambda_{0}\right)_{al}\left(\Sigma_{0}\Lambda_{0}\right)_{aa}^{-1}+\left(\Sigma_{0}\Lambda_{0}\right)_{la}\left(\Sigma_{0}\Lambda_{0}\right)_{ll}^{-1}&\left(\Sigma_{0}\Lambda_{0}\right)_{kl}\left(\Sigma_{0}\Lambda_{0}\right)_{kk}^{-1}+\left(\Sigma_{0}\Lambda_{0}\right)_{lk}\left(\Sigma_{0}\Lambda_{0}\right)_{ll}^{-1}\end{array}\right].$
For all $\left(i,j\right)\in\mathfrak{E}_{G}^{\prime}$, s.t. $\left\{i,j\right\}\notin\left\{\left\{a,k\right\},\left\{a,l\right\},\left\{k,l\right\}\right\}$, it’s clear that
$B_{G_{0}}^{\left[\omega,\sigma_{ij}\right]}=\left(\Sigma_{0}\Lambda_{0}\right)_{ij}\left(\Sigma_{0}\Lambda_{0}\right)_{ii}^{-1}+\left(\Sigma_{0}\Lambda_{0}\right)_{ji}\left(\Sigma_{0}\Lambda_{0}\right)_{jj}^{-1}=0.$
And
$B_{G_{0}}^{\left[\lambda_{ka},\sigma_{ij}\right]}=\delta_{ai}\left[\left(\Sigma_{0}\Lambda_{0}\right)_{kj}\left(\Sigma_{0}\Lambda_{0}\right)_{aa}-\left(\Sigma_{0}\Lambda_{0}\right)_{ka}\left(\Sigma_{0}\Lambda_{0}\right)_{aj}\right]+\delta_{aj}\left[\left(\Sigma_{0}\Lambda_{0}\right)_{ki}\left(\Sigma_{0}\Lambda_{0}\right)_{aa}-\left(\Sigma_{0}\Lambda_{0}\right)_{ka}\left(\Sigma_{0}\Lambda_{0}\right)_{ai}\right].$
When $a=i$, we know that $j\neq a,k,l$, so $\left(\Sigma_{0}\Lambda_{0}\right)_{kj}=\left(\Sigma_{0}\Lambda_{0}\right)_{aj}=0$, thus $\delta_{ai}\left[\left(\Sigma_{0}\Lambda_{0}\right)_{kj}\left(\Sigma_{0}\Lambda_{0}\right)_{aa}-\left(\Sigma_{0}\Lambda_{0}\right)_{ka}\left(\Sigma_{0}\Lambda_{0}\right)_{aj}\right]=0$. Similarly, $\delta_{aj}\left[\left(\Sigma_{0}\Lambda_{0}\right)_{ki}\left(\Sigma_{0}\Lambda_{0}\right)_{aa}-\left(\Sigma_{0}\Lambda_{0}\right)_{ka}\left(\Sigma_{0}\Lambda_{0}\right)_{ai}\right]=0$. Therefore, $B_{G_{0}}^{\left[\lambda_{ka},\sigma_{ij}\right]}=0$, i.e.
$B=\mathbf{0}.$
Use a similar argument, we have
$C=\mathbf{0}.$
Therefore, $B_{G_{0}}$ is in fact a block diagonal matrix:
$B_{G_{0}}=\left[\begin{array}[]{cc}A&\mathbf{0}\\ \mathbf{0}&D\end{array}\right].$
Note that $\det\left(D\right)$ is again a QOP. Let $\Lambda_{0}^{\prime}=2I_{n}$, then $D$ with $\Lambda_{0}=\Lambda_{0}^{\prime}$, denoted as $D^{\prime}$ is diagonal, and all the elements on the diagonal are non-zero. So $\det\left(D^{\prime}\right)\neq 0$. Therefore, the set of values of $\Lambda_{0}$ such that $\det\left(D\right)=0$ has Lebesgue measure zero, i.e. $\det\left(D\right)\neq 0$, generically. On the other hand, let $\Lambda_{0}$ satisfies:
$\lambda_{ij}^{0}=\left\{\begin{array}[]{lcl}2&,&i=j\mbox{ or }\left(i,j\right)=\left(k,a\right)\mbox{ or }\left(i,j\right)=\left(a,l\right)\\ 0&,&\mbox{otherwise}\end{array}\right$.,
then $\det\left(A\right)\neq 0$. This implies that $\det(A)\neq 0$ generically. Because the union of finitely many Lebesgue measure zero sets has Lebesgue measure zero, we conclude that $\det\left(B_{G_{0}}\right)=\det\left(A\right)\det\left(D\right)\neq 0$ generically.
Case 2.
Let $G_{0}=\left(V,\mathfrak{E}_{G_{0}}\right)$ where
$\mathfrak{E}_{G_{0}}=\left\{\left(a,k\right),\left(a,l\right)\right\},$
and
$\lambda_{ij}^{0}=\left\{\begin{array}[]{lcl}2&,&i=j\mbox{ or }\left(i,j\right)=\left(a,k\right)\mbox{ or }\left(i,j\right)=\left(a,l\right)\\ 0&,&\mbox{otherwise}\end{array}\right.$.
Then using similar arguments as in Case 1, we can prove that $\det\left(B_{G_{0}}\right)\neq 0$.
Combining all the arguments above, we can conclude that the set of values of $\Lambda$ such that $\det\left(B_{G}\right)=0$ has Lebesgue measure zero. ∎
Proof of case 11 in Theorem 5.
This is a direct result of Lemmas 9 and $1$. ∎
In conclusion, for graphs without multi-edges, the rank of the Jacobian matrix (the dimension of the model) is calculable.
As explained in Section 2.2, models with different dimensions are generically identifiable. With Theorem 5, the dimensions are known for a subset of models that satisfies Assumption 1. One interpretation of this result is that for a family of models $\left\{\mathbf{M}_{k}\right\}_{k=1}^{K}$ that satisfy Assumption 1 and the condition $n_{r}<n_{c}^{\prime}$, they are generically identifiable if the number of edges in the corresponding graphs is different, i.e. the number of edges is identifiable. See the next section for a summary and illustration of all the results above.
5.3 Summary and illustration
In this section, we first provide a summary of all the identifiability results (Table 1), in which conditions for two stationary VAR(1) models $\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ to satisfy the identifiability criteria or not and corresponding examples are listed. In the table, $\mathfrak{MC}_{i}$ is the set of maximal classes of the model $\mathbf{M}_{i}$.
Table 1: A summary of identifiability results with examples
Conclusion Conditions References Examples
$\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ are identifiable $G_{1}$ and $G_{2}$ do not contain multi-edges, ${\operatorname{dim}}\left(\mathbf{M}_{1}\right)={\operatorname{dim}}\left(\mathbf{M}_{2}\right)$, and $\mathfrak{MC}_{1}\neq\mathfrak{MC}_{2}$ Theorem 3 $G_{1}$
$G_{2}$
${\operatorname{dim}}\left(\mathbf{M}_{1}\right)\neq{\operatorname{dim}}\left(\mathbf{M}_{2}\right)$ Theorem 5 $G_{1}$
$G_{2}$
$G_{1}$ or $G_{2}$ contains multi-edges, and conditions in Theorem 4 are satisfied Theorem 4 Figure 4
$\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ do not satisfy the identifiability criteria in this paper ${\operatorname{dim}}\left(\mathbf{M}_{1}\right)={\operatorname{dim}}\left(\mathbf{M}_{2}\right)$ and $\mathfrak{MC}_{1}=\mathfrak{MC}_{2}$ - $G_{1}$
$G_{2}$
$G_{1}$ or $G_{2}$ contains multi-edges and conditions in Theorem 4 are not satisfied - $G_{1}$
$G_{2}$ is any graph
Note that although we present cases where the identifiability criteria are not satisfied, this does not necessarily mean the models are unidentifiable. These graphs fall into two categories: (i) those containing multi-edges, the dimensions of which cannot be determined using the results in this paper (final row, Table 1); and (ii) those that cannot be distinguished by model dimension or maximal classes (second-to-last row, Table 1). The latter case could suggest a certain degree of homogeneity.
We end this section with an example of a family of graphs that are identifiable from each other.
Example 10.
The family of graphs in Figure 5 are identifiable.
(a) $G_{1}$
(b) $G_{2}$
(c) $G_{3}$
(d) $G_{4}$
Figure 5: A family of identifiable graphs
The relevant properties of the graphs are listed below:
Graph Set of maximal classes Dimension
$G_{1}$ $\left\{\{1,2\},\{1,3\},\{1,4\}\right\}$ 77
$G_{2}$ $\left\{\{1,3\},\{2,3,4\}\right\}$ 88
$G_{3}$ $\left\{\{1,2\},\{2,3,4\}\right\}$ 88
$G_{4}$ $\left\{\{1,2,3\},\{3,4\}\right\}$ Unknown
By the dimensions, $G_{1}$ is identifiable from $G_{2}$ and $G_{3}$. Moreover, $G_{2}$ and $G_{3}$ are identifiable because the maximal classes are different. Finally, $G_{4}$ is identifiable from $G_{1}$, $G_{2}$ and $G_{3}$ because the sets of maximal class do not overlap.
6 Illustrations of ecological networks
This section presents possible applications of the identifiability results introduced in Section 5.
6.1 Bipartite graphs with prior classification
Recall that bipartite graphs refer to a type of graph whose nodes can be divided into two groups, and all edges are directed from one group to the other.
Proposition 10.
Any directed bipartite graphs are (generically) identifiable if the nodes are primarily classified, all edges are directed from one part to the other, and the direction is known.
Example 11.
Figure 6 is an example of a hierarchical ecological network.
Figure 6: Example of hierarchical ecological networks ([11])
The maximal classes are:
$\left\{\{1,6,7\},\{2,6,7,8,9\},\{3,6,7,8,9,10\},\{4,6\},\{5,6\}\right\}.$
If we already know that species 151-5 are resources, and species 6106-10 are consumers, then each maximal class contains exactly one resource and all the consumers it feeds, thus the whole graph is identifiable.
6.2 Bipartite graphs without prior classification
Generally, for a directed bipartite graph, whose edges are directed from one part to the other, and the direction is known, if we do not know the classification of the nodes, the whole graph is not identifiable. However, maximal classes can help to recover at least part of the graph. As shown in the following example.
Example 12.
Consider again the network shown in Figure 6, if we do not know who are the resources, or who are the consumers, then we will not be able to recover the whole graph. But we know that each maximal class contains exactly one resource and all the consumers it feeds.
6.3 Resilience of the network
Maximal classes can sometimes indicate the resilience of the network.
Proposition 11.
If two maximal classes are disjoint, the nodes from one maximal class are completely unrelated to the ones from the other.
Example 13.
Figure 7 shows two examples of ecological networks.
Figure 7: Example of hierarchical ecological networks ([11])
The Maximal classes of $A$ are
$\left\{\{1,2,3\},\{4,5,6\}\right\},$
and the maximal classes of $B$ are
$\left\{\{1,2,3,6\},\{3,4,5,6\}\right\}.$
In this case, $A$ is more resilient than $B$, because the only two maximal classes of $A$ are disjoint, and hence nodes 131-3 do not affect nodes 464-6 at all. In particular, if we take node $3$ away from the system, then in $A$, only nodes $1$ and $2$ will be affected, while in $B$, all the other nodes will be affected.
Acknowledgments
I would like to express my deepest gratitude to my PhD supervisors: Stéphane Robin, Viet Chi Tran and Elisa Thébault. I am especially thankful to Stéphane Robin and Viet Chi Tran, for their unwavering support, guidance, and encouragement throughout the research journey, as well as for the passionate and intellectually stimulating spark they brought to our countless discussions, without which this research would not have been possible. I am also deeply grateful to Elisa Thébault for her valuable contribution in bringing an ecological perspective to this work, which enriched it tremendously. This research benefited from the support of the Chair Modélisation Mathématiques et Biodiversité VEOLIA – École polytechnique-MNHN – FX.
Appendix A Algebraic dimension
This appendix introduces formally the algebraic dimension with respect to Zariski topology in the framework of algebraic geometry ([9]), and explains why the dimension can be defined as the rank of the Jacobian matrix at a generic point.
First, we introduce the notion of variety, on which the algebraic dimension is defined.
Definition 11.
Let $\mathbb{R}\left[X\right]$ be the set of polynomials of $n$ variables over $\mathbb{R}$, where $X\in\mathbb{R}^{n}$, and $S\subseteq\mathbb{R}\left[X\right]$ be a subset. The variety defined by $S$ is the set:
$\mathcal{V}(S):=\left\{X\in\mathbb{R}^{n}\mid f(X)=0,\forall f\in S\right\}.$
The variety $V(S)$ is also called the zero set of $S$. A variety of algebraic geometry is not always a manifold in differential geometry because it could contain singularities (e.g. sharp points). Conversely, varieties without singular points fall into the domain of differential geometry as a smooth manifold.
Now, we define the algebraic dimension of a variety.
Definition 12.
Let $V$ be a variety. Define the dimension of $V$ (denoted as ${\operatorname{dim}}(V)$) to be the supremum of all integers $n$ such that there exists a chain $V_{0}\subsetneq V_{1}\subsetneq\cdots\subsetneq V_{n}$ of distinct subvarieties of $V$.
This definition corresponds to Zariski topology, in which the closed sets are defined as the zero sets of any set of polynomials.
Example 14.
The dimension of $\mathbb{R}^{1}$ is $1$, ([9, Chapter 1]). Indeed, the only subvarieties of $\mathbb{R}^{1}$ are the whole space and single points.
The dimension of non-variety sets is simply the dimension of its algebraic closure, i.e. the smallest variety containing the set. For example, $\mathbb{R^{+}}$ is not a variety. Its dimension is $1$ because its algebraic closure is $\mathbb{R}^{1}$.
Our model $\mathbf{M}_{G}$ is also not a variety, the dimension would refer to the dimension of its algebraic closure. This dimension is calculable by Gröbner bases ([18, Section 3.4]), yet computationally demanding. Instead, we use an equivalent definition for the dimension - the rank of the Jacobian matrix of the parametrization map at a generic point. This is true because the parametrization map is rational (proof in Appendix B) ([18, Chapter 16, Th.16.1.7]).
Appendix B Proof of the fact that the parametrization map is rational
Proof.
Recall that the parametrization map is:
$\phi_{G}:M_{G}\times\mathbb{R}^{+}$ $\rightarrow M_{n}\left(\mathbb{R}\right)$
$\left(\Lambda,\omega\right)$ $\mapsto\Sigma,\mbox{ s.t. }\mbox{vec}\left(\Sigma\right)=\left(I_{n}-\Lambda^{T}\otimes\Lambda^{T}\right)^{-1}\mbox{vec}\left(\omega I_{n}\right).$
It’s sufficient to prove that for all $i\in\left[n^{2}\right]$, the $i^{th}$ coordinate function of $\phi_{G}$:
$\phi_{i}:M_{G}^{B}\times\mathbb{R}^{+}$ $\rightarrow M_{n}\left(\mathbb{R}\right)$
$\left(\Lambda,\omega\right)$ $\mapsto\mbox{vec}\left(\Sigma\right)_{i}$
is a quotient of polynomials (QOPs), i.e. $\exists f_{i},g_{i}\in\mathbb{R}\left[\lambda_{11},\lambda_{12},\cdots,\lambda_{nn},\omega\right],s.t.\phi_{i}=f_{i}/g_{i}$, where $\mbox{vec}\left(\Sigma\right)_{i}$ is the $i^{th}$ element of $\mbox{vec}\left(\Sigma\right)$.
For all $i\in\left[n^{2}\right]$,
$\phi_{i}\left(\Lambda,\omega\right)$ $=\mbox{vec}\left(\Sigma\right)_{i}$
$=\sum_{j=1}^{n^{2}}\left(I_{n^{2}}-\Lambda^{T}\otimes\Lambda^{T}\right)^{-1}_{ij}\mbox{vec}\left(\omega I_{n}\right)_{j}$
$=\omega\sum_{j=1}^{n^{2}}\left(I_{n^{2}}-\Lambda^{T}\otimes\Lambda^{T}\right)^{-1}_{ij}\mbox{vec}\left(I_{n}\right)_{j}$
$=\omega\sum_{k=1}^{n}\left(I_{n^{2}}-\Lambda^{T}\otimes\Lambda^{T}\right)^{-1}_{i,\left(\left(k-1\right)n+k\right)}.$
Since the finite sum of QOPs is still a QOP, it’s sufficient to prove that $\left(I_{n^{2}}-\Lambda^{T}\otimes\Lambda^{T}\right)^{-1}$ is a matrix whose elements are QOPs.
By definition,
$\left(I_{n^{2}}-\Lambda^{T}\otimes\Lambda^{T}\right)^{-1}=\frac{1}{\det\left(I_{n^{2}}-\Lambda^{T}\otimes\Lambda^{T}\right)}\mbox{adj}\left(I_{n^{2}}-\Lambda^{T}\otimes\Lambda^{T}\right),$
where $\mbox{adj}\left(\cdot\right)$ is the adjugate matrix. Since $\det\left(\cdot\right)$ is a polynomial, and $\mbox{adj}\left(\cdot\right)$ is a matrix whose elements are polynomials, $\left(I_{n^{2}}-\Lambda^{T}\otimes\Lambda^{T}\right)^{-1}$ is a matrix whose elements are QOPs. ∎
Appendix C Proof of Lemma 1
Proof.
Recall from (2):
$\Sigma=\Lambda^{T}\Sigma\Lambda+\omega I_{n},$
and for all $i,j\in[n]$,
$\left(\Lambda^{T}\Sigma\Lambda\right)_{ij}=\sum_{k,l=1}^{n}\lambda_{li}\sigma_{lk}\lambda_{kj}=\sum_{k,l=1}^{n}\sigma_{kl}\lambda_{ki}\lambda_{lj}.$
Let
$\delta_{ij}=\left\{\begin{array}[]{lcr}1&\mbox{if}&i=j\\ 0&\mbox{if}&i\neq j\end{array}\right.,\qquad\forall i,j\in[n].$
For all $a,b,i,j\in[n]$,
$\begin{split}\frac{\partial\sigma_{ij}}{\partial\lambda_{ab}}&=\frac{\partial\left(\Lambda^{T}\Sigma\Lambda\right)_{ij}}{\partial\lambda_{ab}}+\frac{\partial\left(\omega I_{n}\right)_{ij}}{\partial\lambda_{ab}}=\frac{\partial\left(\Lambda^{T}\Sigma\Lambda\right)_{ij}}{\partial\lambda_{ab}}\\ &=\sum_{k,l=1}^{n}\frac{\partial\sigma_{kl}}{\partial\lambda_{ab}}\lambda_{ki}\lambda_{lj}+\sum_{k,l=1}^{n}\sigma_{kl}\frac{\partial\left(\lambda_{ki}\lambda_{lj}\right)}{\partial\lambda_{ab}}\\ &=\sum_{k,l=1}^{n}\frac{\partial\sigma_{kl}}{\partial\lambda_{ab}}\lambda_{ki}\lambda_{lj}+2\delta_{ij}\delta_{bi}\sum_{l=1}^{n}\sigma_{al}\lambda_{lb}+\left(1-\delta_{ij}\right)\left(\delta_{bi}\sum_{l=1}^{n}\sigma_{al}\lambda_{lj}+\delta_{bj}\sum_{l=1}^{n}\sigma_{al}\lambda_{li}\right),\\ \frac{\partial\sigma_{ij}}{\partial\omega}&=\frac{\partial\left(\Lambda^{T}\Sigma\Lambda\right)_{ij}}{\partial\omega}+\frac{\partial\left(\omega I_{n}\right)_{ij}}{\partial\omega}=\sum_{k,l=1}^{n}\frac{\partial\sigma_{kl}}{\partial\omega}\lambda_{ki}\lambda_{lj}+\delta_{ij}.\end{split}$ (6)
Let $\overline{\mathbf{J}}_{\lambda_{ab},\sigma_{ij}}=\partial\sigma_{ij}/\partial\lambda_{ab}$ and $\overline{\mathbf{J}}_{\omega,\sigma_{ij}}=\partial\sigma_{ij}/\partial\omega$. Then, using (6),
$\overline{\mathbf{J}}_{\lambda_{ab},\sigma_{ij}}$ $=\sum_{k,l=1}^{n}\overline{\mathbf{J}}_{\lambda_{ab},\sigma_{kl}}A_{\lambda_{kl},\sigma_{ij}}+B_{\lambda_{ab},\sigma_{ij}},$
$\overline{\mathbf{J}}_{\omega,\sigma_{ij}}$ $=\sum_{k,l=1}^{n}\overline{\mathbf{J}}_{\omega,\sigma_{kl}}A_{\lambda_{kl},\sigma_{ij}}+B_{\omega,\sigma_{ij}},$
where
$A_{\lambda_{ab},\sigma_{ij}}$ $=\lambda_{ai}\lambda_{bj},$
$B_{\lambda_{ab},\sigma_{ij}}$ $=2\delta_{ij}\delta_{bi}\sum_{l=1}^{n}\sigma_{al}\lambda_{lb}+\left(1-\delta_{ij}\right)\left(\delta_{bi}\sum_{l=1}^{n}\sigma_{al}\lambda_{lj}+\delta_{bj}\sum_{l=1}^{n}\sigma_{al}\lambda_{li}\right)$
$=2\delta_{ij}\delta_{bi}\left(\Sigma\Lambda\right)_{ab}+\left(1-\delta_{ij}\right)\left(\delta_{bi}\left(\Sigma\Lambda\right)_{aj}+\delta_{bj}\left(\Sigma\Lambda\right)_{ai}\right),$
$B_{\omega,\sigma_{ij}}$ $=\delta_{ij},$
If we order the rows and columns of $\overline{\mathbf{J}}$ such that the rows correspond to $\lambda_{11}$$\lambda_{12}$$\lambda_{1n}$$\lambda_{n1}$$\lambda_{n2}$$\lambda_{nn}$, and $\omega$ respectively, and the columns correspond to $\sigma_{11}$$\sigma_{12}$$\sigma_{1n}$$\sigma_{n1}$$\sigma_{n2}$$\sigma_{nn}$, then we are able to organize AA and $B$ into the following matrix forms:
$A$ $=\Lambda\otimes\Lambda,$
$B$ $=\left[\begin{array}[]{ccccccccccccccc}2\left(\Sigma\Lambda\right)_{11}&\left(\Sigma\Lambda\right)_{12}&\cdots&\left(\Sigma\Lambda\right)_{1n}&\left(\Sigma\Lambda\right)_{12}&0&\cdots&0&&&&\left(\Sigma\Lambda\right)_{1n}&0&\cdots&0\\ 0&\left(\Sigma\Lambda\right)_{11}&\cdots&0&\left(\Sigma\Lambda\right)_{11}&2\left(\Sigma\Lambda\right)_{12}&\cdots&\left(\Sigma\Lambda\right)_{1n}&&&&0&\left(\Sigma\Lambda\right)_{1n}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots&&\cdots&&\vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\left(\Sigma\Lambda\right)_{11}&0&0&\cdots&\left(\Sigma\Lambda\right)_{12}&&&&\left(\Sigma\Lambda\right)_{11}&\left(\Sigma\Lambda\right)_{12}&\cdots&2\left(\Sigma\Lambda\right)_{1n}\\ &&&&&&&&&&&&&&\\ 2\left(\Sigma\Lambda\right)_{21}&\left(\Sigma\Lambda\right)_{22}&\cdots&\left(\Sigma\Lambda\right)_{2n}&\left(\Sigma\Lambda\right)_{22}&0&\cdots&0&&&&\left(\Sigma\Lambda\right)_{2n}&0&\cdots&0\\ 0&\left(\Sigma\Lambda\right)_{21}&\cdots&0&\left(\Sigma\Lambda\right)_{21}&2\left(\Sigma\Lambda\right)_{22}&\cdots&\left(\Sigma\Lambda\right)_{2n}&&&&0&\left(\Sigma\Lambda\right)_{2n}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots&&\cdots&&\vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\left(\Sigma\Lambda\right)_{21}&0&0&\cdots&\left(\Sigma\Lambda\right)_{22}&&&&\left(\Sigma\Lambda\right)_{21}&\left(\Sigma\Lambda\right)_{22}&\cdots&2\left(\Sigma\Lambda\right)_{2n}\\ &&&&&&&&&&&&&&\\ &&&&&&&&&&&&&&\\ &&\vdots&&&&\vdots&&&&&&&\vdots&\\ &&&&&&&&&&&&&&\\ &&&&&&&&&&&&&&\\ 2\left(\Sigma\Lambda\right)_{n1}&\left(\Sigma\Lambda\right)_{n2}&\cdots&\left(\Sigma\Lambda\right)_{nn}&\left(\Sigma\Lambda\right)_{n2}&0&\cdots&0&&&&\left(\Sigma\Lambda\right)_{nn}&0&\cdots&0\\ 0&\left(\Sigma\Lambda\right)_{n1}&\cdots&0&\left(\Sigma\Lambda\right)_{n1}&2\left(\Sigma\Lambda\right)_{n2}&\cdots&\left(\Sigma\Lambda\right)_{nn}&&&&0&\left(\Sigma\Lambda\right)_{nn}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots&&\cdots&&\vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\left(\Sigma\Lambda\right)_{n1}&0&0&\cdots&\left(\Sigma\Lambda\right)_{n2}&&&&\left(\Sigma\Lambda\right)_{n1}&\left(\Sigma\Lambda\right)_{n2}&\cdots&2\left(\Sigma\Lambda\right)_{nn}\\ &&&&&&&&&&&&&&\\ 1&0&\cdots&0&0&1&\cdots&0&&\cdots&&0&0&\cdots&1\end{array}\right]$
$=\left[\begin{array}[]{c}\left(\Sigma\Lambda\right)\otimes I_{n}+\left(\left(\Sigma\Lambda\right)\otimes I_{n}\right)\mathbf{P}\\ \mbox{vec}\left(I_{n^{2}}\right)^{T}\end{array}\right]$
$=\left[\begin{array}[]{c}\left(\Sigma\Lambda\otimes I_{n}\right)\left(I_{n^{2}}+\mathbf{P}\right)\\ \mbox{vec}\left(I_{n^{2}}\right)^{T}\end{array}\right],$
and
$\overline{\mathbf{J}}=\overline{\mathbf{J}}A+B.$
Appendix D Proof of Lemma 2
Proof.
Recall that $\overline{\mathbf{J}}_{G}$ has $E_{G}+1$ rows, which correspond to $\lambda_{ij}$, where $i,j\in\left[n\right]$ and $\left(i,j\right)\in\mathfrak{E}_{G}$ and $\omega$. By definition, $\overline{\mathbf{J}}$ has extra rows comparing to $\overline{\mathbf{J}}_{G}$, and in particular, rows that correspond to $\lambda_{st}$, where $s,t\in\left[n\right]$ and $\left(s,t\right)\notin\mathfrak{E}_{G}$. Therefore the first step of the projection $\psi_{G}$ is to remove these extra rows.
It’s clear that for any $i,j,a,b,s,t\in\left[n\right]$, where $\left(i,j\right)\in\mathfrak{E}_{G}$ and $\left(s,t\right)\notin\mathfrak{E}_{G}$,
$\left(\frac{\partial\sigma_{ab}}{\partial\lambda_{ij}}\right)|_{\lambda_{st}=0}=\frac{\partial\left(\sigma_{ab}|_{\lambda_{st}=0}\right)}{\partial\lambda_{ij}},\mbox{ and }\left(\frac{\partial\sigma_{ab}}{\partial\omega}\right)|_{\lambda_{st}=0}=\frac{\partial\left(\sigma_{ab}|_{\lambda_{st}=0}\right)}{\partial\omega}.$
These equations show that elements in $\overline{\mathbf{J}}_{G}$ are exactly the corresponding elements in $\overline{\mathbf{J}}$ setting $\lambda_{st}$ to 0 for all $s,t\in\left[n\right]$ s.t. $\left(s,t\right)\notin\mathfrak{E}_{G}$. ∎
Appendix E Proof of Lemma 3
Proof.
Denote the interaction matrix and covariance matrix for complete graphs with $n$ nodes as $\Lambda_{c}$ and $\Sigma_{c}$ respectively. We know from Lemma 1 that the extended Jacobian matrix for the complete graph satisfies:
$\overline{\mathbf{J}}=\overline{\mathbf{J}}\left(\Lambda_{c}\otimes\Lambda_{c}\right)+B,$ (7)
where
$B=\left[\begin{array}[]{c}\left(\Sigma_{c}\Lambda_{c}\otimes I_{n}\right)\left(I_{n^{2}}+\mathbf{P}\right)\\ \mbox{vec}\left(I_{n^{2}}\right)^{T}.\end{array}\right].$
Apply the projection $\psi_{G}$ to both sides of the equation, we have
$\overline{\mathbf{J}}_{G}=\overline{\mathbf{J}}_{G}\left(\Lambda\otimes\Lambda\right)+\psi_{G}(B),$
where
$B=\left[\begin{array}[]{c}\left(\Sigma\Lambda\otimes I_{n}\right)\left(I_{n^{2}}+\mathbf{P}\right)\\ \mbox{vec}\left(I_{n^{2}}\right)^{T}.\end{array}\right],$
by the definition of $\psi_{G}$. ∎
Appendix F Proof of Lemma 4
Proposition 12.
For any graph $G$, each node belongs to at least one maximal class.
Proof.
Consider any node in the graph, if it has in-degree 0, then it is a source node, and belongs to the maximal class with itself being the source node. If it has a positive in-degree, then it belongs to the same maximal class as its parent node. ∎
Proof of Lemma 4.
$\Rightarrow$” By Definition 9, $i$ and $j$ belong to the same maximal class if and only if there exists a source node $k\in V$, s.t. they are both reachable by $k$. If $k=i$ or $k=j$, then the first condition is satisfied. On the other hand, if $k\neq i,j$, then let $l=k$, thus the second condition is satisfied.
$\Leftarrow$” First, assume the first condition is satisfied, and w.l.o.g, suppose there exists a directed path from $i$ to $j$. $i$ must belong to at least one maximal class in $G$, denoted as $\mathcal{MC}_{[k]}$ with a source $k$. Then there exists a directed path from $k$ to $i$, thus there also exists a path: $k\rightsquigarrow i\rightsquigarrow j$, i.e. $j\in\mathcal{MC}_{[k]}$.
Next, assume the second condition is satisfied. The result is proved using similar arguments as before and the fact that $l$ must belong to at least one maximal class in $G$. ∎
Appendix G Proof of Lemma 5
Proof.
First, we prove that $\mu_{G}\left(M_{G}\backslash M^{P}\right)=0$, i.e., $\Lambda\in M^{P}$ generically. For all $i,j\in V$, s.t. there exists a directed path in $G$ from $i$ to $j$, it’s clear that for all $k\in\mathbb{N}$, $\lambda_{ij}^{[k]}$ is a polynomial of entries of $\Lambda$. Consider $\Lambda_{0}\in M_{G}$ s.t. $\lambda_{st}>0$ for all $(s,t)\in\mathfrak{E}_{G}$, we know that there exists $K\in\mathbb{N}$ s.t. $\forall k>K$, $\lambda_{0,ij}^{[k]}\neq 0$. Therefore, $\forall k>K$, $\lambda_{ij}^{[k]}\neq 0$ generically, i.e. $\mu_{G}\left(M_{G}\backslash M^{P}\right)=0$.
Next, we prove that $\mu_{G}\left(M_{G}\backslash M^{S}\right)=0$. Define:
$M^{S_{1}}$ $:=\left\{\Lambda\in M_{n}\left(\mathbb{R}\right)\mid\forall i,j\in V,\sum_{k=0}^{+\infty}\sum_{a=1}^{n}\lambda_{ai}^{\left[k\right]}\lambda_{aj}^{\left[k\right]}=0\Leftrightarrow\forall k\in\mathbb{N},a\in\left[n\right],\lambda_{ai}^{\left[k\right]}\lambda_{aj}^{\left[k\right]}=0\right\},$
$M^{S_{2}}$ $:=\left\{\Lambda\in M_{n}\left(\mathbb{R}\right)\mid\forall i,j\in V,\sum_{s=1}^{n}\sum_{k=0}^{+\infty}\sum_{l=1}^{n}\lambda_{li}^{\left[k\right]}\lambda_{ls}^{\left[k\right]}\lambda_{sj}=0\Leftrightarrow\forall k\in\mathbb{N},s,l\in\left[n\right],\lambda_{li}^{\left[k\right]}\lambda_{ls}^{\left[k\right]}\lambda_{sj}=0\right\}.$
By definition, $M^{S}=M^{S_{1}}\cap M^{S_{2}}$. Here, we present only the proof of $\mu_{G}\left(M_{G}\backslash M^{S_{1}}\right)=0$, the proof of $\mu_{G}\left(M_{G}\backslash M^{S_{2}}\right)=0$ is omitted because it uses the exact same technique as the other one.
Denote
$h_{ij}=\sum_{k=0}^{+\infty}\sum_{a=1}^{n}\lambda_{ai}^{\left[k\right]}\lambda_{aj}^{\left[k\right]}.$
$h_{ij}$ is a function defined on $M_{G}$, i.e. a function of entries of $\lambda_{st}$ where $(s,t)\in\mathfrak{E}_{G}$. Assume that for some $i,j\in V$, there exists $\Lambda_{0}\in M_{G}$, s.t. there exist $k_{0}\in\mathbb{N}$ and $a_{0}\in[n]$ s.t. $\lambda_{0,a_{0}i}^{[k_{0}]}\lambda_{0,a_{0}j}^{[k_{0}]}>0$, and $h_{ij}\left(\Lambda_{0}\right)=0$. Denote $c_{i}=\lambda_{0,a_{0}i}^{[k_{0}]}$, and $c_{j}=\lambda_{0,a_{0}j}^{[k_{0}]}$. Note that $c_{i}$ and $c_{j}$ are constants in $\mathbb{R}$. Moreover, define another function $h_{ij}|_{k_{0},a_{0}}$ on $M_{G}$ as $h_{ij}$ with $\lambda_{a_{0}i}^{[k_{0}]}=c_{i}$ and $\lambda_{a_{0}j}^{[k_{0}]}=c_{j}$ fixed as constants. Then $h_{ij}|_{k_{0},a_{0}}$ is an analytic function of the entries of $\Lambda$. Consider again $\Lambda^{\prime}\in M_{G}$ s.t. $\lambda_{st}^{\prime}>0$ for all $(s,t)\in\mathfrak{E}_{G}$, then by definition, $h_{ij}|_{k_{0},a_{0}}>0$. Therefore, $h_{ij}|_{k_{0},a_{0}}$ does not constantly equal zero on $M_{G}$. In this case,
$\mu_{G}\left(\left\{\Lambda\in M_{G}\mid h_{ij}|_{k_{0},a_{0}}=0\right\}\right)=0,$
because the set of roots of a non-zero real analytic function has Lebesgue measure zero ([6, Section 3.1]). This implies that
$\mu_{G}\left(\left\{\Lambda\in M_{G}\mid\lambda_{a_{0}i}^{[k_{0}]}\lambda_{a_{0}j}^{[k_{0}]}>0,\mbox{ and }h_{ij}=0\right\}\right)=0.$
Therefore,
$\mu_{G}\left(M_{G}\backslash M^{S_{1}}\right)$ $=\mu_{G}\left(\left\{\Lambda\in M_{G}\mid\exists i,j\in Vs.t.\exists k_{0}\in\mathbb{N},a_{0}\in[n],\lambda_{a_{0}i}^{[k_{0}]}\lambda_{a_{0}j}^{[k_{0}]}\neq 0,\mbox{ and }h_{ij}=0\right\}\right)$
$=\mu_{G}\left(\left\{\Lambda\in M_{G}\mid\exists i,j\in Vs.t.\exists k_{0}\in\mathbb{N},a_{0}\in[n],\lambda_{a_{0}i}^{[k_{0}]}\lambda_{a_{0}j}^{[k_{0}]}>0,\mbox{ and }h_{ij}=0\right\}\right)$
$=\mu_{G}\left(\bigcup_{i,j\in V}\bigcup_{k_{0}\in\mathbb{N}}\bigcup_{a_{0}\in[n]}\left\{\Lambda\in M_{G}\mid\lambda_{a_{0}i}^{[k_{0}]}\lambda_{a_{0}j}^{[k_{0}]}>0,\mbox{ and }h_{ij}=0\right\}\right)$
$\leq\sum_{i,j\in V}\sum_{k_{0}\in\mathbb{N}}\sum_{a_{0}\in[n]}\mu_{G}\left(\left\{\Lambda\in M_{G}\mid\lambda_{a_{0}i}^{[k_{0}]}\lambda_{a_{0}j}^{[k_{0}]}>0,\mbox{ and }h_{ij}=0\right\}\right)$
$=0.$
The inequality is true because $\mathbb{N}$ is countable ([3, Theorem 10.2]). ∎
Appendix H Proof of Lemma 7
Proof.
Recall that $\overline{\mathbf{J}}_{G}$ is $\mathbf{J}_{G}$ with additional columns that coincide with existing columns of $\mathbf{J}_{G}$. Therefore,
$rank\left(\mathbf{J}_{G}\right)=rank\left(\overline{\mathbf{J}}_{G}\right).$
Since
$\overline{\mathbf{J}}_{G}=\psi_{G}\left(B\right)\left(I_{n^{2}}-\Lambda\otimes\Lambda\right)^{-1},$
$rank\left(\overline{\mathbf{J}}_{G}\right)=rank\left(\psi_{G}\left(B\right)\right)$. Therefore,
$rank\left(\mathbf{J}_{G}\right)=rank\left(\psi_{G}\left(B\right)\right).$
Appendix I Proof of Lemma 9
Proof.
By Lemma 7, $\mbox{rank}\left(\mathbf{J}_{G}\right)=\mbox{rank}\left(\psi_{G}(B)\right)$, and therefore it is sufficient to prove that $\mbox{rank}\left(\psi_{G}(B)\right)=n_{r}$.
Consider a $\left(E_{G}+1\right)\times\left(E_{G}+1\right)$ submatrix of $\psi_{G}\left(B\right)$, where the columns correspond to the set
$\left\{\sigma_{ij}\mid i,j\in\left[n\right],(i,j)\mbox{ or }(j,i)\in\mathfrak{E}_{G}\right\}\cup\left\{\sigma_{kl}\right\},$
where $\left\{k,l\right\}\in\mathfrak{C}_{G}^{mc}$. Note that $\mathfrak{C}_{G}^{mc}\neq\emptyset$ by Lemma 8. After reordering the rows and columns, the submatrix, denoted as $\psi_{G}\left(B\right)^{\left[E_{G}+1\right]}$ is
$\psi_{G}\left(B\right)^{\left[E_{G}+1\right]}=\left[\begin{array}[]{ccccc}2\left(\Sigma\Lambda\right)_{11}&0&\cdots&0&\left(\Sigma\Lambda\right)_{1b}\delta_{a1}+\left(\Sigma\Lambda\right)_{1a}\delta_{b1}\\ 0&2\left(\Sigma\Lambda\right)_{22}&\cdots&0&\left(\Sigma\Lambda\right)_{2b}\delta_{a2}+\left(\Sigma\Lambda\right)_{2a}\delta_{b2}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&2\left(\Sigma\Lambda\right)_{nn}&\left(\Sigma\Lambda\right)_{nb}\delta_{an}+\left(\Sigma\Lambda\right)_{na}\delta_{bn}\\ &&&&\\ 2\left(\Sigma\Lambda\right)_{i1}\delta_{j1}&2\left(\Sigma\Lambda\right)_{i2}\delta_{j2}&\cdots&2\left(\Sigma\Lambda\right)_{in}\delta_{jn}&\left(\Sigma\Lambda\right)_{ib}\delta_{ja}+\left(\Sigma\Lambda\right)_{ia}\delta_{jb}\\ &&&&\\ 1&1&\cdots&1&0\end{array}\right].$
Since $n_{r}\leq n_{c}^{\prime}$, it’s sufficient to prove $\psi_{G}\left(B\right)^{\left[E_{G}+1\right]}$ is full rank.
Consider $\psi_{G}\left(B\right)^{\left[E_{G}+1\right]}$ as a block matrix such that
$\psi_{G}\left(B\right)^{\left[E_{G}+1\right]}=\left[\begin{array}[]{cc}A&B\\ C&D\end{array}\right],$
where
$A=\left[\begin{array}[]{cccc}2\left(\Sigma\Lambda\right)_{11}&0&\cdots&0\\ 0&2\left(\Sigma\Lambda\right)_{22}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&2\left(\Sigma\Lambda\right)_{nn}\end{array}\right].$
It’s clear that $A$ is invertible, since $\left(\Sigma\Lambda\right)_{ii}\neq 0$ for all $i\in\left[n\right]$. Apply the Guttman rank additivity formula (see [8]), we have
$rank\left(\psi_{G}\left(B\right)^{\left[E_{G}+1\right]}\right)=rank\left(A\right)+rank\left(D-CA^{-1}B\right).$
Therefore, $\psi_{G}\left(B\right)^{\left[E_{G}+1\right]}$ is full rank if and only if $D-CA^{-1}B$ is full rank.
By definition, for any $\left(i,j\right)\in\mathfrak{E}_{G}^{\prime}$ and $\left(a,b\right)\in S^{\prime}$,
$\left(CA^{-1}B\right)^{\left[\lambda_{ij},\sigma_{ab}\right]}$
$=\left[\begin{array}[]{cccc}2\left(\Sigma\Lambda\right)_{i1}\delta_{j1}&2\left(\Sigma\Lambda\right)_{i2}\delta_{j2}&\cdots&2\left(\Sigma\Lambda\right)_{in}\delta_{jn}\end{array}\right]\left[\begin{array}[]{cccc}2^{-1}\left(\Sigma\Lambda\right)_{11}^{-1}&0&\cdots&0\\ 0&2^{-1}\left(\Sigma\Lambda\right)_{22}^{-1}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&2^{-1}\left(\Sigma\Lambda\right)_{nn}^{-1}\end{array}\right]\left[\begin{array}[]{c}\left(\Sigma\Lambda\right)_{1b}\delta_{a1}+\left(\Sigma\Lambda\right)_{1a}\delta_{b1}\\ \left(\Sigma\Lambda\right)_{2b}\delta_{a2}+\left(\Sigma\Lambda\right)_{2a}\delta_{b2}\\ \vdots\\ \left(\Sigma\Lambda\right)_{nb}\delta_{an}+\left(\Sigma\Lambda\right)_{na}\delta_{bn}\end{array}\right]$
$=\left[\begin{array}[]{ccccc}0&\cdots&\left(\Sigma\Lambda\right)_{ij}\left(\Sigma\Lambda\right)_{jj}^{-1}&\cdots&0\end{array}\right]\left[\begin{array}[]{c}\left(\Sigma\Lambda\right)_{1b}\delta_{a1}+\left(\Sigma\Lambda\right)_{1a}\delta_{b1}\\ \left(\Sigma\Lambda\right)_{2b}\delta_{a2}+\left(\Sigma\Lambda\right)_{2a}\delta_{b2}\\ \vdots\\ \left(\Sigma\Lambda\right)_{nb}\delta_{an}+\left(\Sigma\Lambda\right)_{na}\delta_{bn}\end{array}\right]$
$=\delta_{ja}\left[\left(\Sigma\Lambda\right)_{ij}\left(\Sigma\Lambda\right)_{jj}^{-1}\left(\Sigma\Lambda\right)_{jb}\right]+\delta_{jb}\left[\left(\Sigma\Lambda\right)_{ij}\left(\Sigma\Lambda\right)_{jj}^{-1}\left(\Sigma\Lambda\right)_{ja}\right],$
and
$\left(CA^{-1}B\right)^{\left[\omega,\sigma_{ab}\right]}$
$=\left[\begin{array}[]{cccc}1&1&\cdots&1\end{array}\right]\left[\begin{array}[]{cccc}2^{-1}\left(\Sigma\Lambda\right)_{11}^{-1}&0&\cdots&0\\ 0&2^{-1}\left(\Sigma\Lambda\right)_{22}^{-1}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&2^{-1}\left(\Sigma\Lambda\right)_{nn}^{-1}\end{array}\right]\left[\begin{array}[]{c}\left(\Sigma\Lambda\right)_{1b}\delta_{a1}+\left(\Sigma\Lambda\right)_{1a}\delta_{b1}\\ \left(\Sigma\Lambda\right)_{2b}\delta_{a2}+\left(\Sigma\Lambda\right)_{2a}\delta_{b2}\\ \vdots\\ \left(\Sigma\Lambda\right)_{nb}\delta_{an}+\left(\Sigma\Lambda\right)_{na}\delta_{bn}\end{array}\right]$
$=\left[\begin{array}[]{cccc}2^{-1}\left(\Sigma\Lambda\right)_{11}^{-1}&2^{-1}\left(\Sigma\Lambda\right)_{22}^{-1}&\cdots&2^{-1}\left(\Sigma\Lambda\right)_{nn}^{-1}\end{array}\right]\left[\begin{array}[]{c}\left(\Sigma\Lambda\right)_{1b}\delta_{a1}+\left(\Sigma\Lambda\right)_{1a}\delta_{b1}\\ \left(\Sigma\Lambda\right)_{2b}\delta_{a2}+\left(\Sigma\Lambda\right)_{2a}\delta_{b2}\\ \vdots\\ \left(\Sigma\Lambda\right)_{nb}\delta_{an}+\left(\Sigma\Lambda\right)_{na}\delta_{bn}\end{array}\right]$
$=2^{-1}\left(\Sigma\Lambda\right)_{ab}\left(\Sigma\Lambda\right)_{aa}^{-1}+2^{-1}\left(\Sigma\Lambda\right)_{ba}\left(\Sigma\Lambda\right)_{bb}^{-1}.$
Thus,
$\left(D-CA^{-1}B\right)^{\left[\lambda_{ij},\sigma_{ab}\right]}$ $=\delta_{ja}\left[\left(\Sigma\Lambda\right)_{ib}-\left(\Sigma\Lambda\right)_{ij}\left(\Sigma\Lambda\right)_{jj}^{-1}\left(\Sigma\Lambda\right)_{jb}\right]+\delta_{jb}\left[\left(\Sigma\Lambda\right)_{ia}-\left(\Sigma\Lambda\right)_{ij}\left(\Sigma\Lambda\right)_{jj}^{-1}\left(\Sigma\Lambda\right)_{ja}\right],$
$\left(D-CA^{-1}B\right)^{\left[\omega,\sigma_{ab}\right]}$ $=-2^{-1}\left(\Sigma\Lambda\right)_{ab}\left(\Sigma\Lambda\right)_{aa}^{-1}-2^{-1}\left(\Sigma\Lambda\right)_{ba}\left(\Sigma\Lambda\right)_{bb}^{-1}.$
Therefore, for $B_{G}$ of Definition 10,
$B_{G}^{\left[\lambda_{ij},\sigma_{ab}\right]}$ $=\left(\Sigma\Lambda\right)_{jj}\left(D-CA^{-1}B\right)^{\left[\lambda_{ij},\sigma_{ab}\right]},$
$B_{G}^{\left[\omega,\sigma_{ab}\right]}$ $=\left(-2\right)\left(D-CA^{-1}B\right)^{\left[\omega,\sigma_{ab}\right]}.$
Note that
$B_{G}=M_{G}\left(D-CA^{-1}B\right),$
where $M_{G}$ is a diagonal matrix whose elements on the diagonal are $\left(\Sigma\Lambda\right)_{jj}$, s.t. $\exists i\in\left[n\right],\left(i,j\right)\in\mathfrak{E}_{G}^{\prime}$ and 2-2. Hence $M_{G}$ is invertible, and
$rank\left(D-CA^{-1}B\right)=rank\left(B_{G}\right).$
Appendix J Proof of Case 2 in Theorem 5
By Lemma 8, we know that under Assumption 1, the condition $n_{r}>n_{c}^{\prime}$ is satisfied if and only if the set $\mathfrak{C}_{G}^{mc}$ is empty, i.e. $C_{G}^{mc}=0$, meaning that for all $i,j\in[n]$, either $(i,j)\mbox{ or }(j,i)\in\mathfrak{E}_{G}$ or $i,j$ do not belong to the same maximal class. In addition, $n_{c}^{\prime}=E_{G}$.
Using similar arguments as before, we define another matrix, $B_{G}^{\prime}$, which is derived from the Jacobian matrix, and prove that the Jacobian matrix $\mathbf{J}_{G}$ is full rank if and only if $B_{G}^{\prime}$ is full rank.
Definition 13.
Let $\mathbf{M}_{G}$ be a stationary VAR(1) model that satisfies $n_{r}>n_{c}^{\prime}$. Define a square matrix $B_{G}^{\prime}$ of size $E_{G}^{\prime}$, where the rows correspond to $\lambda_{ij}$, s.t. $i,j\in[n]$ and $(i,j)\in\mathfrak{E}_{G}^{\prime}$, and the columns correspond to $\sigma_{ab}$ s.t. $a,b\in[n]$ and $(a,b)$ or $(b,a)\in\mathfrak{E}_{G}^{\prime}$ as follows
$B_{G}^{\prime\left[\lambda_{ij},\sigma_{ab}\right]}=\delta_{ja}\left[\left(\Sigma\Lambda\right)_{ib}\left(\Sigma\Lambda\right)_{jj}-\left(\Sigma\Lambda\right)_{ij}\left(\Sigma\Lambda\right)_{jb}\right]+\delta_{jb}\left[\left(\Sigma\Lambda\right)_{ia}\left(\Sigma\Lambda\right)_{jj}-\left(\Sigma\Lambda\right)_{ij}\left(\Sigma\Lambda\right)_{ja}\right].$
Note that $B_{G}^{\prime}$ is $B_{G}$ excluding the last row that corresponds to $\omega$.
Lemma 12.
Let $\mathbf{M}_{G}$ be a stationary VAR(1) model that satisfies the condition $n_{r}>n_{c}^{\prime}$. Then $\mbox{rank}\left(\psi_{G}\left(B\right)\right)=n_{c}^{\prime}$ (i.e. $\mbox{rank}\left(\mathbf{J}_{G}\right)=n_{c}^{\prime}$) if $B_{G}^{\prime}$ is full rank, i.e. $\mbox{rank}(B_{G}^{\prime})=E_{G}^{\prime}$.
Proof.
Consider a $E_{G}\times E_{G}$ submatrix of $\psi_{G}(B)$, where the rows correspond to $\lambda_{ij}$ s.t. $i,j\in[n]$ and $(i,j)\in\mathfrak{E}_{G}$, and the columns correspond to $\sigma_{ab}$ s.t. $\in[n], aba\leq b$, and $(a,b)$ or $(b,a)\in\mathfrak{E}_{G}$. In fact, this submatrix contains all distinct non-zero columns of $\psi_{G}(B)$. Therefore, if this submatrix is full rank, then $\mbox{rank}\left(\psi_{G}\left(B\right)\right)=n_{c}^{\prime}=E_{G}$.
Use the same technique as in Lemma 9, where we conder $\psi_{G}\left(B\right)$ as a block matrix and apply the Guttman rank additivity formula, we know that
$\mbox{rank}(\psi_{G}(B))=E_{G}\ \Leftrightarrow\ \mbox{rank}(B_{G}^{\prime})=E_{G}^{\prime}.$
Denote a subset $\overline{M_{G}^{B}}$ of $M_{G}$:
$\overline{M_{G}^{B}}:=\left\{\Lambda\in M_{G}\mid\det(B_{G}^{\prime})=0\right\}.$
Lemma 13.
Let $\mathbf{M}_{G}$ be a stationary VAR(1) model that satisfies Assumption 1 and $n_{r}>n_{c}^{\prime}$, then
$\mu_{G}\left(\overline{M_{G}^{B}}\right)=0.$
Proof.
Use similar arguments as in Lemma 11, $\det(B_{G}^{\prime})$ is a rational function of the entries of $\Lambda$ and $\omega$. Therefore, it is sufficient to find a $\Lambda_{0}\in M_{G}$ s.t. $\det(\overline{B_{G_{0}}})\neq 0$, where $\overline{B_{G_{0}}}$ is $B_{G}^{\prime}$ with $\Lambda=\Lambda_{0}$. Let
$\Lambda_{0}=2I_{n},$
then $\overline{B_{G_{0}}}$ is diagonal with non-zero diagonal entries. Therefore $\det(B_{G}^{\prime})\neq 0$. ∎